### v1

# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. While hyperscalers are collectively failing to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations, due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced, by adjusting the fraction of requests served by different service quality tiers. We show, that by adapting the the quality of responses with respect to local carbon intensity, we can achieve additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a multi-horizon optimization, that reaches close-to-optimal carbon savings under realistic conditions, and can dynamically adapt the service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
Carbon-aware computing, quality of service, LLM inference, generative AI, service-oriented computing
submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v1. Under review.
1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which is contributing directly to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming (particularly for cloud gaming and virtual reality), the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
Figure 1. The carbon intensity of electricity can vary significantly over time. Exploiting these temporal variations can reduce the carbon footprint of energy-intensive services.
Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi). While cloud providers generally report Scope 1 emissions (direct emissions from company-owned or controlled sources, such as on-site fuel combustion), reporting on Scope 2 emissions (i.e. electricity consumed by data centers) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI). Moreover, upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), which include all indirect emissions that occur across the value chain. These include emissions from the use of of cloud services, which will affect not only service providers but also customers of energy-intensive services.
The concept of carbon-aware computing has emerged as a response to these challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency (i.e., reducing demand) by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. However, existing carbon-aware approaches primarily focus on determining the ideal time and location of executing flexible batch processing tasks (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024; Sukprasert et al., 2024b). Services, on the other hand, usually require permanent availability and have little to no delay-tolerance, which is why carbon efficiency in services has been mainly explored in the context of geo-spatial load shifting (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). Yet, many services cannot simply be distributed across regions due to latency constraints (i.e., proximity of servers and users), data availability, privacy regulations, security concerns, and—most importantly—infrastructure limitations. For instance, generative AI inference typically relies on high-performance GPUs or other accelerators, that are available only at specific locations.
Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to mere energy efficiency.
In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only reduce energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric, that describes the proportion of requests that are served by different service quality tiers. For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021). Similarly, in LLM inference, adjusting the number of returned tokens is another way to directly impact the energy demand of requests (Li et al., 2024b).
Adjusting service quality is common practice for avoiding overload (Meza et al., 2023; Park et al., 2024), but has also recently been proposed as a strategy for improving service sustainability (Vitali, 2022). Figure 2 illustrates, how carbon-aware QoR adaptions can reduce emissions by an additional 5-10 % in many regions—beyond the savings related to reduced energy demand.
Contributions
Towards more carbon-efficient cloud services, we make the following contributions in this paper:
  1. (1)
    We formalize the problem of minimizing carbon emissions under QoR constraints through optimized resource provisioning and load balancing between service quality tiers.
  2. (2)
    We model emissions for data center operators and cloud tenants (Scope 2 and 3 reporting) and show that both perspectives are equivalent while optimizing for energy and carbon efficiency.
  3. (3)
    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can dynamically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
  4. (4)
    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service across two request traces in ten different regions.
All results from this analysis are fully reproducible. Code and data will be made publicly available upon acceptance.
2. Background and Motivation
We review the current state of GHG reporting for data centers and cloud services, and motivate our work by outlining anticipated regulatory changes.
2.1. Energy-Intensive Cloud Services
The growing computational demand of generative AI is a major factor for the expansion of data centers worldwide (S&P Globl Market Intelligence, 2023). In particular, the rapid adoption of large language models (LLMs) requires immense processing power, not only for initial training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024). As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase electricity demand in the future (Li et al., 2024a).
Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands. Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024). Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024). Additionally, scientific simulations—including climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), and computational fields like astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022)—are growing in complexity and increasingly integrate AI inference services.
2.2. Carbon Footprint of Cloud Providers
Despite previous commitments to achieve fully carbon-free or even ‘carbon-negative’ operations by 2030, Google has reported a 48 % increase in GHG emissions in the last five years (Google, 2024), Microsoft has announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI. The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
All big hyperscalers have announced or introduced carbon reporting tools for their customers (Google Cloud, 2023; Microsoft, 2023; Amazon Web Services, 2023; Dixit and Tse, 2024). However, they are expected to significantly underreport emissions, due to opaque and inconsistent reporting methodologies (O’Brien, 2024). We anticipate an increasing focus on regulatory frameworks that will require cloud providers to disclose their carbon emissions more comprehensively.
2.3. Growing Pressure from GHG Reporting
The GHG Protocol (Protocol, 2001) is a globally recognized framework that provides standards for organizations to report their GHG emissions. It forms the basis for many national and international climate programs, such as the Carbon Disclosure Project (CDP) (Project, 2024) and Science Based Targets initiative (SBTi) (initiative, SBTi) and categorizes emissions into three scopes:
  • Scope 1: Direct GHG emissions that occur from sources that are owned or controlled by the company.
  • Scope 2: Indirect GHG emissions from the generation of purchased electricity consumed by the company.
  • Scope 3: All other indirect emissions that occur across the value chain, such as from suppliers or services.
For the IT sector, Scope 2 and 3 emissions are the most relevant categories. Their reporting is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024). While, the disclosure of Scope 3 emissions stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023). Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019). From 2026 on, this will also include non-EU companies with significant operations in the EU.
Although current legislation around the world mostly addresses Scope 1 and 2 emissions, the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018). Therefore, also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
2.4. Implications for Energy-Intensive Cloud Services
Figure 3. Service providers either operate their own infrastructure or rent from cloud providers. Either way, electricity usage will be reflected in GHG reporting, as operational carbon emissions are passed on to cloud tenants. Data center Scope 3 emissions are often reffered to as embodied emissions.
When reporting emissions from the perspective of providers of energy-intensive services, we must distinguish between two cases, as illustrated in Figure 3:
  1. (1)
    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity purchases, i.e. Scope 2 emissions.
  2. (2)
    They rent infrastructure from cloud providers, in which case the data center’s Scope 2 and 3 emissions will be passed down the value chain, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023).
While we will address the two cases in our problem formalization, both are affected by stricter GHG reporting, as outlined earlier. Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost. Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
2.5. Assumptions on Future Regulatory Policies
As the effectiveness of yearly and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions. Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024; Sukprasert et al., 2024b; Radovanovic et al., 2022). ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source. Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time. Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework. In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
3. Carbon-Aware QoR Adaptation
Offering diverse service quality tiers for different user groups is common practice in the operation of cloud services. For example, content delivery networks provide faster and more reliable service to higher-tier users, while AI platforms like OpenAI and Gemini reserve access to larger and more advanced models for paying customers. Furthermore, service providers already degrade service quality during periods of high load to prevent system overload (Park et al., 2024; Meza et al., 2023). Such measures often target best-effort users: For instance, free-tier users of LLM inference services can experience highly increased latencies at certain times, to not endanger the QoS of subscribed users.
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency. For instance, service quality tiers for LLMs can range from entirely different models (e.g. number of parameters), quantization (many models existin in variants from 4bit integers to 32bit floating point numbers), as well as number of returned tokens (i.e. we generate shorter replies for some users), which can have a significant impact on the carbon footprint of inference (Li et al., 2024b).
In this work, we investigate the effect of adapting service QoR over time for improved carbon efficiency within a single region, i.e., without geo-distributed load balancing.
3.1. Problem Setting
We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time periods
(1) $\mathcal{T}_{i}=[i\Delta,(i+1)\Delta],\quad i=0,1,\dots,I-1,$
where $I=T/\Delta$ is the total number of intervals, i.e. $i=0$ corresponds to the first interval and $i=I-1$ to the last. For the remainder of this paper, we consider $T=1~{}\text{year}$ and $\Delta=1$ hour.
We define the set of user groups as $\mathcal{U}$, set of machine types $\mathcal{M}$ and set of service quality tiers $\mathcal{Q}$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests which are allocated to different service quality tiers ($A^{i}$). The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Note that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines. Table 1 shows an overview of all input and decision variables.
Table 1. Description of symbols. Small letters denote the elements of the corresponding, capitalized matrix.
Input variables
$\mathcal{U}$ Set of user groups $u$
$\mathcal{M}$ Set of machine types $m$
$\mathcal{Q}$ Set of service quality tiers $q$
$\gamma$ Validity period length
$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $\mathcal{T}_{i}$.
$C^{i}\in\mathbb{R}$ Carbon intensity during $\mathcal{T}_{i}$.
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\left(k_{m,q}\right)$ that machine $m$ can serve at quality $q$.
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\left(p^{i}_{m,q}\right)$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
$C_{m}^{\text{emb}}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
Decision variables
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\left(d^{i}_{m,q}\right)$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\left(a^{i}_{u,q}\right)$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
3.2. Quality of Responses
We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers. In our analysis, we focus on a single user group, such as best-effort users, whose requests can be fulfilled by two distinct tiers that differ in both response quality and energy demand. In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
In the following, we introduce several metrics to establish a more general definition of QoR.
Service level indicators
Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), we define the service level indicator $\text{SLI}(\alpha,\omega)\in[0,1]^{\left|U\right|\times\left|Q\right|}$. It describes the observed distribution of requests across different service tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
(2) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
Service level objectives
To describe the minimum QoR requirements that a service must satisfy, users of this framework can define a service-level objective $\text{SLO}\in[0,1]^{\left|U\right|\times\left|Q\right|}$. The SLO indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$.
Similarly, we define a service level upper-bound $\text{SLU}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality. For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
[…]
$\text{SLO}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{SLU}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
describes that requests by premium users must always be served by Tier 2, while 60-100 % of requests best-effort users are served by Tier 1 and up to 40 % by Tier 2.
As all requests for each user group have to be served by some quality
(3) $\sum_{q\in\mathcal{Q}}\text{slo}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\text{slu}_{u,q}=1\quad\forall u\in\mathcal{U}.$
General definition of QoR
We define the service’s QoR within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between SLI (Equation 2), SLO and SLU:
(4) $\text{QoR}(\alpha,\omega)=1-\min_{u\in\mathcal{U},q\in\mathcal{Q}}\left(\frac{\left|\text{sli}_{u,q}(\alpha,\omega)-\text{slo}_{u,q}\right|}{\left|\text{slu}_{u,q}-\text{slo}_{u,q}\right|}\right),$
meaning that the largest deviation from SLU will determine the final QoR. Following the above example, an
$\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$
would result in a $\text{QoR}(\alpha,\omega)=0.75$.
Validity periods
As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$. These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
Figure 4 illustrates how varying the length of validity periods influences the distribution of hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
Figure 4. Increasing validity periods allow for more flexibility in adjusting the proportion of requests served by Tier 1 and Tier 2 based on carbon intensity.
3.3. Power Model
To determine the operational carbon emissions from powering servers, we first define the power usage $p^{i}_{m,q}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$. We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
Data center operator perspective
[…]
(5) $p^{i}_{m,q}=\text{PUE}\cdot\left(p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(\text{util}^{i}_{q}\right)^{n}\right),$
where
(6) $\text{util}^{i}_{q}=\frac{\sum_{u\in\mathcal{U}}a^{i}_{u,q}}{\sum_{m\in\mathcal{M}}\left(d^{i}_{m,q}\cdot k_{m,q}\right)}.$
Cloud tenant perspective
Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions. Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change. Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor. We assume that current methodologies attribute a constant factor $P_{m}^{\text{ estimate}}$ to each machine $m$ which does not change over time or depend on the executed workload or utilization of the machine:
(7) $p^{i}_{m,q}=p_{m}^{\text{estimate}}$
3.4. Carbon Emission Model
We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ in respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied carbon $C_{m}^{\text{ emb}}$ (in gCO2)
(8) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\cdot\left(p^{i}_{m,q}\cdot C^{i}+C_{m}^{\text{emb}}\right)$
Depending on methodology, embodied emissions might not always be reported with respect to the usage of infrastructure, in which case $C_{m}^{\text{emb}}=0$.
3.5. Optimization Problem
We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies the $\text{QoR}_{\text{target}}$ across validity periods.
(9) $\min\quad$ $\sum_{i=\alpha}^{\omega}E^{i}$
(10) s.t. $\sum_{q\in\mathcal{Q}}a^{i}_{u,q}=r^{i}_{u}$ $\forall i,\forall u$
(11) $\sum_{u\in\mathcal{U}}a^{i}_{u,q}\leq\sum_{m\in\mathcal{M}}d^{i}_{m,q}\cdot k_{m,q}$ $\forall i,\forall q$
(12) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
where
  • All requests served (10) ensures that all requests during each interval $\mathcal{T}_{i}$ are attributed to a quality tier.
  • Sufficient resources (11) ensures that the provided machines have sufficient capacity to serve the requests allocated to each quality tier.
  • Sufficient QoR (12) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
Note, that this model assumes perfect knowledge of the future (in particular perfect forecasts for all $C^{i}$ and $R^{i}$) and can therefore not be solved in practice. However, we use it as an upper bound for assessing potential and as the foundation of the online proposed in Section 4.
Complexity
From a cloud tenant perspective (Scope 3), the optimization problem is an NP-hard mixed-integer linear program (MILP), as the objective and all constraints are linear, but involve the discrete decision variables $D^{i}$.
From the perspective of data center operators, the Scope 2 power model from Equation (5) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve. However, during optimization, we can relax Equation (5) to the same form as Equation (7), by assuming $p^{i}_{m,q}=p_{m,q}^{\text{max}}$, as long as $n\leq 1$, i.e. the power model is concave. If the power law relationship is strictly convex, i.e. $n>1$, it becomes more efficient to deploy multiple machines at lower utilization levels $util^{i}_{q}$ rather than fully utilizing fewer machines. However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007). Hence, for $n\leq 1$, the power models for Scope 2 and Scope 3 reporting are equivalent during optimization, as they result in the same decisions for optimal $A^{i}$ and $D^{i}$.
Lastly, when values in $D^{i}$ are sufficiently big (i.e. in large-scale scenarios with many active machines), the optimal result can be approximated by relaxing $D^{i}$ to continuous variables. This turns the problem into a simple linear program that is solvable in polynomial time.
4. Online Approach
For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$ as time progresses. Its performance is limited by two factors:
  1. (1)
    the errors of request and carbon intensity forecasts
  2. (2)
    the required approximations for solving the NP-hard optimization problem
Yet, our experiments show that even under realistic forecasts and strict time limits for solving the MILP, we can achieve close-to-optimal emission savings (Section 5.4).
4.1. Forecasting
We denote forecasts for request volume and carbon intensity as $\widehat{R}^{i}$ and $\widehat{C}^{i}$, respectively. There are a variety of different approaches commonly for long-term forecasts with multiple seasonalities, like Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011). As the electricity mix of regions does usually not change abruptly, long-term forecasts for ACI can be estimated reasonably well by fitting daily, weekly, and yearly seasonalities. However, especially in regions that heavily rely on weather-dependent renewable power production, additionally utilizing short-term forecasts (typically up to 4 days) provides significantly better accuracy (Maji et al., 2023; Li et al., 2024c). For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
4.2. Multi-Horizon Optimization
Based on the optimization problem introduced in Section 3.5, we introduce a multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$. The approach is described in Algorithm 1.
Algorithm 1 Multi-Horizon Optimization
1:for $\alpha\leftarrow 0$ to $I-1$ do
2:    # Long-term optimization
3:   if $\alpha\equiv 0\pmod{\tau}$ then
4:      $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
5:      $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\sum_{i=\alpha}^{I-1}E^{i}$
6:    # Short-term optimization
7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}}\sum_{i=\alpha}^{I-1}E^{i}$
8:    # Progress in time
[…]
10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality. To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
  1. (1)
    A long-term optimization, executed every $\tau$ intervals (for example, every 24 hours), solves the MILP for the remainder of the year. Any $D^{i}$ and $A^{i}$ in the past are fixed, i.e. we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval. In this step, longer execution times and less optimal solutions are acceptable.
  2. (2)
    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^1] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (11) constraint.
After optimization, the determined deployment $D^{\alpha}$ is provisioned for the current interval. As short-term load and carbon intensity predictions are usually very precise, and the provisioning of new nodes can take significant time (Microsoft reports 6-8 minutes for creating a new LLM inference instance (Stojkovic et al., 2024)) we assume that there are usually no rapid auto-scaling decisions within an interval. After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
4.3. Dynamic QoR Adaptation for Best-Effort Users
The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, another relevant question is: What are the expected yearly emissions from this user group, and can we keep them within a predefined budget?
In Section 5.5, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected yearly emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year. Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:
(13) $\max\quad$ $\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
[…]
(14) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
In the multi-horizon optimization (Algorithm 1), we replace the long-term optimization in Line 5 with this objective:
(15) $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\max_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\left(\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=0}^{I-1}\right)$
Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate the emissions from all intervals in Equation (14). This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$. Past values in $D^{i}$ and $A^{i}$ are fixed.
5. Experiments
We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the entire year 2023 using two real-life request traces and in ten different regions.
All experiments were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job. We use Gurobi (Gurobi Optimization, LLC, 2024) for solving the optimization problems.
5.1. Experimental Setup
As average carbon intensity is usually reported hourly, we consider $\Delta=1$ hour.
Scenario
[…]
We model machine performance using recent benchmarks of vLLM on a EC2 p4d.24xlarge for inference, which state a throughput of 11.57 requests per second for LLaMA 3.1 8B and 5.05 requests per second for LLaMA 3.1 70B (vLLM Team, [n. d.]).
Request traces
We simulate realistic request patterns using two hourly request traces from the Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), as shown in Figure 5. In particular, we use
  • Wiki (en): the English Wikipedia with $63.4e6\pm 0.8e$ 6 requests per hour and spikes of up to $616.4e$ 6 requests per hour. In our scenario, this results in 80-200 active machines throughout the year.
  • Wiki (de): the German Wikipedia with $54.7e5\pm 2.8e$ 5 requests per hour, with spikes of up to $517.6e$ 5 requests per hour. This results in only 10-25 active machines, emphasizing the importance of discrete scheduling decisions.
Figure 5. The request traces differ notably, as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
Carbon intensity regions
We analyze the potential in ten different regions:
  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
  • In Australia, we consider Queensland (AUS-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalize all time zones to align patterns with the time of day in the request traces. Temporal variability of hourly carbon intensities are shown in Figure 1.
5.2. Impact of QoR on Energy Demand
We first analyze the absolute yearly emissions generated by our different scenarios without carbon-aware QoR adaption. For this, we solve Equation (9) for both request traces, across all 10 regions, $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$, and $\gamma=1$ hour. The results reflect emission savings driven purely by energy demand, i.e., by serving more or less requests on Llama 3.1 70B. We optimize until the objective is within 0.1 % of its optimal solution or after 10 minutes. All results are within 0.8 % of the optimal solution.
Figure 6 shows the overall emissions for Wiki (en). As expected, we observe very large differences across different regions. For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally. Serving all requests with the 80B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source. As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$. Results are equivalent for Wiki (de), but smaller by a factor of around 7.5.
Figure 6. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ for all 10 regions.
Figure 7. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains due to increased flexibility.
5.3. Upper-Bound Potential of Carbon-Aware QoR Adaptation
[…]
Carbon Savings Beyond Energy Efficiency
As define scenarios with validity periods of $\gamma=1$ hour as baselines. For these scenarios, emissions are directly correlated with energy demand as there is no flexibility for carbon-aware QoR adaptation. Figure 7 reports the relative and absolute additional savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods with respect to these baselines.
Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we can see more notable potential of 5-8 % in some regions. The characteristics of carbon intensity data significantly influence the expected gains. For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany experiences more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal impact in California, whereas in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity, which limits the potential for carbon-aware approaches.
Impact of $\text{QoR}_{\text{target}}$
Figure 8. Additional relative savings for different $\text{QoR}_{\text{target}}$ in Germany and California.
Figure 8 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$. When $\text{QoR}_{\text{target}}$ is 0, the optimizer will serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings. Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B. Values around $\text{QoR}_{\text{target}}=0.5$ offer the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Figure 9. Additional carbon savings under realistic conditions. Our online approach maintains additional savings within one percentage point across most experiments.
Impact of validity periods
(a) $\beta=\{1,4,8,24\}$ hours for smaller validity periods.
(b) $\beta=\{1,3,7\}$ days for larger validity periods.
Figure 10. Considering all time intervals of length $\beta$ throughout the year, what proportion of these intervals have a QoR below the threshold? We analyze the CDF for Wiki (en) in Germany for $\text{QoR}_{\text{target}}=0.5$.
While increasing $\gamma$ enhances potential for carbon savings, it also raises the risk of prolonged periods with low QoR. Service providers can adjust the range of QoR using SLO and SLG to ensure that a certain fraction of requests can always be served from a specific quality tier. However, limiting the fraction of requests available for QoR adaptation lowers its potential impact.
Figure 10 illustrates the probability of prolonged periods with low QoR under different validity period lengths. To to provide an example of how to interpret the results: Figure 10(a) shows that when optimizing for a $\text{QoR}_{\text{target}}=0.5$ with validity periods of $\gamma=8$ hours, no 8-hour interval ($\beta$) has a QoR below 0.5, while all intervals maintain a QoR of at least 0.5. However, around 10 % of all 4-hour intervals fall below a QoR of 0.25.
We observe, that for validity periods of $\gamma=3$ months, users can experience several consecutive days of low QoR. Consequently, we conclude that very large validity periods of more than a month are not practical in real-world scenarios.
5.4. Potential under Realistic Conditions
We evaluate the practicability of carbon-aware QoR adaptation using our proposed online optimization under real forecasts and near-optimal MILP solutions.
Setup
We simulate request forecasts using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and yearly seasonalities. In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year. The predictions exhibit realistic errors: For Wiki (en), 24-hour forecasts have an average MAPE of $13.9\pm 8.4$, while for Wiki (de), the error is $32.1\pm 15.3$.
For long-term carbon intensity forecasts, we predict the entire year of 2023 as of January 1st using Prophet, fitted on three years of historical data. We simulate short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region (see Appendix A). Short-term forecasts are updated daily at midnight. We interrupt long-term optimizations after 30s and short-term optimizations after 10s, to demonstrate performance under near-optimal solutions.
Furthermore, we limit this analysis to daily, weekly, and monthly validity periods, as Section 5.3 showed that validity periods shorter than a day offer minimal potential for carbon savings, while periods longer than a month can lead to extended durations of low QoR.
Results
Figure 9 shows the performance gap of our online optimization compared to the theoretical upper bound established in Section 5.3. Across all load traces and regions, the resulting emissions savings are generally within one percentage point of the optimal solution. In regions with already limited upper-bound potential, the expected savings become more marginal. However, in most regions, our approach consistently maintains additional savings of 5–10 %.
Within the stated time limits, 95 % of all long-term optimizations achieve an optimization gap $<2.1\,\%$, while 95 % of all short-term optimizations reach a gap $<0.3\,\%$.
5.5. Outlook of Yearly Emissions
(a) Yearly emission outlook at three discrete points in time.
(b) Yearly emission outlook, performed weekly for $\text{QoR}_{\text{target}}=0.5$.
Figure 11. Projected yearly emissions (blue) based on long-term forecasts for different $\text{QoR}_{\text{target}}$. The interval shows the range between best- and worst-case forecasts of $\widehat{R}$, while red indicates emissions already incurred.
Long-term optimizations become computationally cheap when accepting near-optimal solutions. This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts $\widehat{R}$, which may represent, for instance, worst-case or best-case scenarios.
Setup
We demonstrate emission outlooks by executing the online approach on the Wiki (en) traces in California. For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively. Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
Results
Figure 11(a) shows the resulting outlooks a the start of the year, after 10 weeks, and after 40 weeks. The outlooks provide a comprehensive picture of the future, based on the currently available information. From the initial outlook (left), service providers may conclude that setting $\text{QoR}_{\text{target}}=0.5$ is a reasonable target to stay within an annual carbon budget of 1000 tCO2. Consequently, we optimize for $\text{QoR}_{\text{target}}=0.5$ for the remainder of the year.
Figure 11(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$. We observe that the upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as actual emissions accumulate.
5.6. Dynamic QoR Adaptation for Best-Effort Users
Figure 12. Daily QoR for California, Poland, and Sweden. Our adaptive approach provides the most consistent QoR throughout the year.
Setup
We illustrate the dynamic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.3 on Wiki (en) in California, Poland, and Sweden to cover a spectrum of differently behaving carbon intensity traces: California exhibits high variability due to many solar deployments, Poland has relatively little variation due to reliance on fossil fuels, and Sweden usually has very low carbon intensity which occasionally spikes when demand needs to be met through cross-border imports or natural gas. For each region, we define the annual budget $B^{\text{annual}}$ so that the ideal $\text{QoR}_{\text{target}}=0.50$.
We compare our approach to two greedy baselines:
  1. (1)
    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves Equation (13). Any remaining or overspent budget is evenly distributed across the remaining time steps.
  2. (2)
    Greedy (weighted) solves Equation (13), too, but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as forecasts are updated, weightings get adjusted.
Results
Figure 12 illustrates the daily QoR of our online approach compared to both greedy baselines over time. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR fluctuates significantly throughout the year. Greedy (weighted) delivers more consistent QoR for both Poland and Sweden, though it starts very conservatively in both regions, leading to near-perfect QoR at the year’s end due to unused budget. In California, where forecasts are less accurate, this baseline overspends early in the year, resulting in days with $\text{QoR}=0$ for most of the year’s end. Our approach also starts conservatively, as the forecast initially overestimates annual emissions (see Figure 11). However, it delivers the most consistent service across all experiments.
6. Discussion
We highlight the future impact of carbon efficiency in energy-intensive services and line out the limitations of this study.
Impact
We want to emphasize the extend of carbon efficiency improvements for cloud services, based on our analysis in Section 4.3. ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our scenarios. Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy. However, and increasing part of the community is advocating to base load shifting decisions based on marginal data (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids. However, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone. Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion. We predict, that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods.
Limitations
In this paper, we provide first insights into the potential of carbon-aware QoR adaptation based on an LLM inference use case. Applying our approach to more complex, real-world scenarios with multiple user groups, service quality tiers, and heterogeneous infrastructure is possible from a modeling perspective, but not yet investigated. Furthermore, we only concentrate on QoR within a single data center and do not distribute load across multiple geo-distributed regions—which is known to yield the largest potential for carbon savings (Sukprasert et al., 2024b). This paper represents the first comprehensive analysis of QoR adaptation for energy and carbon efficiency, and sets the stage for more detailed explorations in future research.
Second, our current modeling only considers the QoR for user groups, rather than individual users. This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven. For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group. We consider addressing fairness aspects with more granular user modeling as as an area for future work.
7. Related Work
We survey related works on carbon-aware computing as well as carbon efficiency in energy-intensive cloud services.
Carbon-aware computing
The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy. Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024; Sukprasert et al., 2024b). However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed datacenters and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016). For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint while respecting latency constraints. CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks. Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity. However, no existing approaches address services that must operate in specific data centers due to data, security, or infrastructure constraints.
Sustainability of cloud services
Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023). Especially the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024).
While service degradation is a common practice to prevent overload (Meza et al., 2023; Park et al., 2024), it has already been proposed to design applications to offer different quality tiers of a service for sustainability (Vitali, 2022). In the context of carbon-aware LLMs, there only exists a single approach, Sprout (Li et al., 2024b), which optimizes an LMMs autoregressive generation process. In particular, the authors propose to return shorter responses during periods of high carbon intensity, as the energy demand of a response directly correlates with the amount of generated tokens. Sprout clearly classifies as an approach which adapts QoR over time and can benefit from our scheduling decisions.
8. Conclusion
As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets. In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency. Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand alone. Moreover, we propose an approach for dynamically adjusting the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
Acknowledgements.
This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. 01IS18025A and 01IS18037A).
Appendix A CarbonCast MAPE
Region Day 1 Day 2 Day 3 Day 4
CISO 8.08 11.19 12.93 13.62
PJM 3.69 4.93 5.87 6.67
ERCOT 9.78 10.93 11.61 12.23
NYISO 6.91 9.06 9.95 10.42
SE 4.29 5.64 6.43 6.74
DE 7.81 10.69 12.80 15.55
PL 3.12 4.14 4.72 5.50
ES 10.12 16.00 19.37 21.12
NL 6.06 7.87 9.08 9.99
AU-QLD 3.93 3.98 4.06 5.87
Table 2. Carbon forecast MAPE over 96 hours for all observed regions as reported in CarbonCast (Maji et al., 2023)

### v2

@@
 The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace.
-While hyperscalers are collectively failing to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions.
+While hyperscalers collectively fail to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions.
 To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing.
-However, such approaches are not applicable to services that require constant availability at specific locations, due to latency, privacy, data, or infrastructure constraints.
-In this paper, we explore how the carbon footprint of energy-intensive services can be reduced, by adjusting the fraction of requests served by different service quality tiers.
-We show, that by adapting the the quality of responses with respect to local carbon intensity, we can achieve additional carbon savings beyond resource and energy efficiency.
-Building on this, we introduce a multi-horizon optimization, that reaches close-to-optimal carbon savings under realistic conditions, and can dynamically adapt the service quality for best-effort users to stay within an annual carbon budget.
+However, such approaches are not applicable to services that require constant availability at specific locations due to latency, privacy, data, or infrastructure constraints.
+In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers.
+We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency.
+Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget.
 Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
-Carbon-aware computing, quality of service, LLM inference, generative AI, service-oriented computing
-submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v1. Under review.
+Sustainable computing, quality of service, LLM inference, green AI
+submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v2. Under review.
 1. Introduction
-Operating modern cloud services can require substantial computing resources and energy, which is contributing directly to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023).
-Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming (particularly for cloud gaming and virtual reality), the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024).
+Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023).
+Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024).
 As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
@@
 Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi).
-While cloud providers generally report Scope 1 emissions (direct emissions from company-owned or controlled sources, such as on-site fuel combustion), reporting on Scope 2 emissions (i.e.
-electricity consumed by data centers) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI).
-Moreover, upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), which include all indirect emissions that occur across the value chain.
-These include emissions from the use of of cloud services, which will affect not only service providers but also customers of energy-intensive services.
-The concept of carbon-aware computing has emerged as a response to these challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency (i.e., reducing demand) by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is.
-However, existing carbon-aware approaches primarily focus on determining the ideal time and location of executing flexible batch processing tasks (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024; Sukprasert et al., 2024b).
-Services, on the other hand, usually require permanent availability and have little to no delay-tolerance, which is why carbon efficiency in services has been mainly explored in the context of geo-spatial load shifting (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013).
-Yet, many services cannot simply be distributed across regions due to latency constraints (i.e., proximity of servers and users), data availability, privacy regulations, security concerns, and—most importantly—infrastructure limitations.
-For instance, generative AI inference typically relies on high-performance GPUs or other accelerators, that are available only at specific locations.
-Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to mere energy efficiency.
-In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only reduce energy demand, but can yield additional carbon savings.
-We define the quality of responses (QoR) as a type of quality of service (QoS) metric, that describes the proportion of requests that are served by different service quality tiers.
+While reporting on Scope 2 emissions (GHG emissions caused by electricity consumption) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI), upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023)—including emissions from the use of cloud services.
+This will affect not only service providers, but also users of energy-intensive services.
+The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is.
+Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a).
+However, for continuously running services, where individual requests have little to no delay-tolerance, existing approaches exclusively focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013).
+However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations.
+For instance, generative AI inference typically relies on high-performance GPUs or specialized accelerators, which are often only available at specific locations.
+Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
+Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
+In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings.
+We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by different service quality tiers.
 For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021).
-Similarly, in LLM inference, adjusting the number of returned tokens is another way to directly impact the energy demand of requests (Li et al., 2024b).
-Adjusting service quality is common practice for avoiding overload (Meza et al., 2023; Park et al., 2024), but has also recently been proposed as a strategy for improving service sustainability (Vitali, 2022).
-Figure 2 illustrates, how carbon-aware QoR adaptions can reduce emissions by an additional 5-10 % in many regions—beyond the savings related to reduced energy demand.
-Contributions
-Towards more carbon-efficient cloud services, we make the following contributions in this paper:
+Our approach minimizes the operational carbon emissions of energy-intensive services by optimizing resource provisioning and load balancing across service quality tiers, while ensuring compliance with service level objectives (SLOs) for different user groups.
+Figure 2 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week.
+Serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0).
+Balancing requests evenly between the two tiers results in intermediate energy usage.
+We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints.
+Based on this, we propose an approach that achieves close-to-optimal savings despite real-life challenges such as forecast inaccuracies and the high computational complexity of the underlying optimization.
+Contributions.  Towards more carbon-efficient cloud services, we make the following contributions:
   1. (1)
@@
   2. (2)
-    We model emissions for data center operators and cloud tenants (Scope 2 and 3 reporting) and show that both perspectives are equivalent while optimizing for energy and carbon efficiency.
+    We model emissions from the perspective of data center operators and cloud tenants (Scope 2 and 3 reporting) and prove that both models lead to equivalent decisions.
   3. (3)
-    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can dynamically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
+    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can automatically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
   4. (4)
-    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service across two request traces in ten different regions.
-All results from this analysis are fully reproducible. Code and data will be made publicly available upon acceptance.
+    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service on eight real and synthetic request traces in ten different regions.
+All experiments in this paper are fully reproducible. All code and data will be made publicly available upon acceptance.
 2. Background and Motivation
-We review the current state of GHG reporting for data centers and cloud services, and motivate our work by outlining anticipated regulatory changes.
-2.1. Energy-Intensive Cloud Services
-The growing computational demand of generative AI is a major factor for the expansion of data centers worldwide (S&P Globl Market Intelligence, 2023).
-In particular, the rapid adoption of large language models (LLMs) requires immense processing power, not only for initial training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024).
-As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase electricity demand in the future (Li et al., 2024a).
-Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands.
-Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024).
-Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024).
-Additionally, scientific simulations—including climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), and computational fields like astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022)—are growing in complexity and increasingly integrate AI inference services.
-2.2. Carbon Footprint of Cloud Providers
-Despite previous commitments to achieve fully carbon-free or even ‘carbon-negative’ operations by 2030, Google has reported a 48 % increase in GHG emissions in the last five years (Google, 2024), Microsoft has announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI.
+Despite previous commitments to achieve fully carbon-free or even “carbon-negative” operations by 2030, Microsoft recently announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), Google reported a 48 % increase in GHG emissions within five years (Google, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI.
 The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
@@
 We anticipate an increasing focus on regulatory frameworks that will require cloud providers to disclose their carbon emissions more comprehensively.
-2.3. Growing Pressure from GHG Reporting
-The GHG Protocol (Protocol, 2001) is a globally recognized framework that provides standards for organizations to report their GHG emissions.
-It forms the basis for many national and international climate programs, such as the Carbon Disclosure Project (CDP) (Project, 2024) and Science Based Targets initiative (SBTi) (initiative, SBTi) and categorizes emissions into three scopes:
-  • Scope 1: Direct GHG emissions that occur from sources that are owned or controlled by the company.
-  • Scope 2: Indirect GHG emissions from the generation of purchased electricity consumed by the company.
-  • Scope 3: All other indirect emissions that occur across the value chain, such as from suppliers or services.
-For the IT sector, Scope 2 and 3 emissions are the most relevant categories.
-Their reporting is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
+Energy-intensive cloud services.
+The rapid adoption of large language models (LLMs) requires immense processing power, not only for training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024).
+As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase future electricity demand (Li et al., 2024a).
+Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands.
+Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024).
+Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024).
+Additionally, scientific simulations such as climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022) are growing in complexity and increasingly integrate AI services.
+Growing pressure from GHG reporting.
+The reporting of Scope 2 emissions, i.e.
+indirect GHG emissions caused by consuming electricity, is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
 In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024).
-While, the disclosure of Scope 3 emissions stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023).
+While, the disclosure of Scope 3 emissions—i.e.
+all indirect emissions that occur across the value chain—stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023).
 Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019).
 From 2026 on, this will also include non-EU companies with significant operations in the EU.
-Although current legislation around the world mostly addresses Scope 1 and 2 emissions, the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018).
-Therefore, also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
-2.4. Implications for Energy-Intensive Cloud Services
-Figure 3.
-Service providers either operate their own infrastructure or rent from cloud providers.
-Either way, electricity usage will be reflected in GHG reporting, as operational carbon emissions are passed on to cloud tenants.
-Data center Scope 3 emissions are often reffered to as embodied emissions.
-When reporting emissions from the perspective of providers of energy-intensive services, we must distinguish between two cases, as illustrated in Figure 3:
+As the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018), also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
+Implications for energy-intensive services.  When reporting emissions from the perspective of cloud service providers, we must distinguish between two cases:
   1. (1)
-    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity purchases, i.e. Scope 2 emissions.
+    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity usage, i.e. Scope 2 emissions.
   2. (2)
-    They rent infrastructure from cloud providers, in which case the data center’s Scope 2 and 3 emissions will be passed down the value chain, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
+    They rent infrastructure from cloud providers, in which case the data center’s emissions will be passed down the value chain als Scope 3 emissions, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
 Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023).
-While we will address the two cases in our problem formalization, both are affected by stricter GHG reporting, as outlined earlier.
+We will address the two cases in our problem formalization, however, both are affected by stricter GHG reporting.
 Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost.
 Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
-2.5. Assumptions on Future Regulatory Policies
-As the effectiveness of yearly and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions.
+Assumptions on future regulatory policies.
+As the effectiveness of annual and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions.
 Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
-We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024; Sukprasert et al., 2024b; Radovanovic et al., 2022).
+We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024b; Radovanovic et al., 2022).
 ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source.
@@
 3. Carbon-Aware QoR Adaptation
-Offering diverse service quality tiers for different user groups is common practice in the operation of cloud services.
+Offering different service quality tiers is common practice in the operation of cloud services.
 For example, content delivery networks provide faster and more reliable service to higher-tier users, while AI platforms like OpenAI and Gemini reserve access to larger and more advanced models for paying customers.
@@
 Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency.
-For instance, service quality tiers for LLMs can range from entirely different models (e.g.
-number of parameters), quantization (many models existin in variants from 4bit integers to 32bit floating point numbers), as well as number of returned tokens (i.e.
-we generate shorter replies for some users), which can have a significant impact on the carbon footprint of inference (Li et al., 2024b).
-In this work, we investigate the effect of adapting service QoR over time for improved carbon efficiency within a single region, i.e., without geo-distributed load balancing.
+For instance, service quality tiers for LLMs can vary based on factors such as the choice of model (e.g., number of parameters), quantization (ranging from 4-bit integers to 32-bit floating-point), and the number of tokens generated (i.e.
+providing shorter replies (Li et al., 2024b)), and can have a significant impact on the carbon footprint of inference.
+In this work, we examine the impact of dynamically adjusting service QoR over time to enhance carbon efficiency within a single region, without relying on geo-distributed load balancing.
 3.1. Problem Setting
-We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time periods
+We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time intervals
 (1) $\mathcal{T}_{i}=[i\Delta,(i+1)\Delta],\quad i=0,1,\dots,I-1,$
@@
 For the remainder of this paper, we consider $T=1~{}\text{year}$ and $\Delta=1$ hour.
-We define the set of user groups as $\mathcal{U}$, set of machine types $\mathcal{M}$ and set of service quality tiers $\mathcal{Q}$.
+We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$ and service quality tiers $\mathcal{Q}$.
 Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
@@
 Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$.
-The deployment $D^{i}$ must be sufficient to meet the number of incoming requests which are allocated to different service quality tiers ($A^{i}$).
+The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$) [^1] Note, that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines.
+Furthermore, when elements in $D^{i}$ are sufficiently large (i.e.
+scenarios with many active machines), an optimal result can be approximated by relaxing $D^{i}$ to continuous variables.
+This turns the optimization problem introduced later into a linear program that can be solved in polynomial time..
 The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval.
-Note that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines.
 Table 1 shows an overview of all input and decision variables.
@@
 $C^{i}\in\mathbb{R}$ Carbon intensity during $\mathcal{T}_{i}$.
-$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\left(k_{m,q}\right)$ that machine $m$ can serve at quality $q$.
-$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\left(p^{i}_{m,q}\right)$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
+$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
+$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
 $C_{m}^{\text{emb}}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
 Decision variables
-$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\left(d^{i}_{m,q}\right)$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
-$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\left(a^{i}_{u,q}\right)$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
+$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
+$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
 3.2. Quality of Responses
 We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers.
-In our analysis, we focus on a single user group, such as best-effort users, whose requests can be fulfilled by two distinct tiers that differ in both response quality and energy demand.
+In our subsequent analysis, we concentrate on a single user group, such as best-effort users, whose requests can be served by two distinct service tiers that differ in both response quality and energy demand.
 In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2.
 When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
-In the following, we introduce several metrics to establish a more general definition of QoR.
-Service level indicators
-Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), we define the service level indicator $\text{SLI}(\alpha,\omega)\in[0,1]^{\left|U\right|\times\left|Q\right|}$.
-It describes the observed distribution of requests across different service tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
-(2) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
-Service level objectives
-To describe the minimum QoR requirements that a service must satisfy, users of this framework can define a service-level objective $\text{SLO}\in[0,1]^{\left|U\right|\times\left|Q\right|}$.
-The SLO indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$.
-Similarly, we define a service level upper-bound $\text{SLU}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality.
+In the following, we introduce several metrics to establish a more general definition of QoR, that holds for multiple user groups and service quality tiers.
+Service level objectives.
+Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), users of this framework can define a lower-bound service level objective $\underline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$, where $\underline{\text{\text{slo}}}_{u,q}$ indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$.
+Similarly, users can define an upper-bound $\overline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality.
 For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
 […]
-$\text{SLO}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{SLU}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
-describes that requests by premium users must always be served by Tier 2, while 60-100 % of requests best-effort users are served by Tier 1 and up to 40 % by Tier 2.
-As all requests for each user group have to be served by some quality
-(3) $\sum_{q\in\mathcal{Q}}\text{slo}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\text{slu}_{u,q}=1\quad\forall u\in\mathcal{U}.$
-General definition of QoR
-We define the service’s QoR within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between SLI (Equation 2), SLO and SLU:
-(4) $\text{QoR}(\alpha,\omega)=1-\min_{u\in\mathcal{U},q\in\mathcal{Q}}\left(\frac{\left|\text{sli}_{u,q}(\alpha,\omega)-\text{slo}_{u,q}\right|}{\left|\text{slu}_{u,q}-\text{slo}_{u,q}\right|}\right),$
-meaning that the largest deviation from SLU will determine the final QoR. Following the above example, an
-$\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$
-would result in a $\text{QoR}(\alpha,\omega)=0.75$.
-Validity periods
+$\underline{\text{SLO}}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{and}\quad\overline{\text{SLO}}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
+describes that premium users’ requests must always be served by Tier 2, while 60–100 % of requests by best-effort users can be served by Tier 1 and 0–40 % by Tier 2.
+All requests per user group must be served by a quality tier:
+(2) $\sum_{q\in\mathcal{Q}}\underline{\text{slo}}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\overline{\text{slo}}_{u,q}=1\quad\forall u\in\mathcal{U}.$
+Service level indicators.
+The service level indicator $\text{SLI}(\alpha,\omega)$ describes the observed distribution of requests across different quality tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
+(3) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
+As the service level is constrained by the SLOs, each
+(4) $\text{sli}_{u,q}(\alpha,\omega)\in\left[\min(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q}),\max(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q})\right].$
+General definition of QoR.
+We define the service’s quality of responses $\text{QoR}(\alpha,\omega)$ within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between $\text{SLI}(\alpha,\omega)$, $\underline{\text{SLO}}$ and $\overline{\text{SLO}}$.
+In particular, for the set of parameters that permit optimization, defined as $\mathcal{V}=\{(u,q)\in\mathcal{U}\times\mathcal{Q}:\overline{\text{slo}}_{u,q}\neq\underline{\text{slo}}_{u,q}\}$, we define QoR as:
+(5) $\text{QoR}(\alpha,\omega)=1-\min_{(u,q)\in\mathcal{V}}\left(\frac{\big{|}\overline{\text{slo}}_{u,q}-\text{sli}_{u,q}(\alpha,\omega)\big{|}}{\big{|}\overline{\text{slo}}_{u,q}-\underline{\text{slo}}_{u,q}\big{|}}\right).$
+Following the above example, an $\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$ would result in a $\text{QoR}(\alpha,\omega)=0.75$.
+Validity periods.
 As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$.
@@
 Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
-Figure 4 illustrates how varying the length of validity periods influences the distribution of hourly QoR over time.
+Figure 3 illustrates how the validity period length influences the hourly QoR over time.
 Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
@@
 3.3. Power Model
-To determine the operational carbon emissions from powering servers, we first define the power usage $p^{i}_{m,q}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$.
+To determine the operational carbon emissions from powering servers, we define the power usage $P^{i}=\bigl{(}p^{i}_{m,q}\bigl{)}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$.
 We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
-Data center operator perspective
 […]
@@
 (7) $\text{util}^{i}_{q}=\frac{\sum_{u\in\mathcal{U}}a^{i}_{u,q}}{\sum_{m\in\mathcal{M}}\left(d^{i}_{m,q}\cdot k_{m,q}\right)}.$
-Cloud tenant perspective
+Cloud tenant perspective.
 Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions.
@@
 Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor.
-We assume that current methodologies attribute a constant factor $P_{m}^{\text{ estimate}}$ to each machine $m$ which does not change over time or depend on the executed workload or utilization of the machine:
-(7) $p^{i}_{m,q}=p_{m}^{\text{estimate}}$
+We assume that cloud providers attribute a constant, load-independent factor $P_{m}^{\text{attr}}$ to each active machine $m$:
+(8) $p^{i}_{m,q}=p_{m}^{\text{attr}}$
 3.4. Carbon Emission Model
-We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ in respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied carbon $C_{m}^{\text{ emb}}$ (in gCO2)
+We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ with respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2):
 (9) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\cdot\left(p^{i}_{m,q}\cdot C^{i}+C_{m}^{\text{emb}}\right)$
-Depending on methodology, embodied emissions might not always be reported with respect to the usage of infrastructure, in which case $C_{m}^{\text{emb}}=0$.
+Depending on methodology, embodied emissions might not always be reported with respect to infrastructure usage, in which case $C_{m}^{\text{emb}}=0$.
 3.5. Optimization Problem
-We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies the $\text{QoR}_{\text{target}}$ across validity periods.
+We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies $\text{QoR}_{\text{target}}$:
 (10) $\min\quad$ $\sum_{i=\alpha}^{\omega}E^{i}$
@@
 (12) $\sum_{u\in\mathcal{U}}a^{i}_{u,q}\leq\sum_{m\in\mathcal{M}}d^{i}_{m,q}\cdot k_{m,q}$ $\forall i,\forall q$
+[…]
 (14) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
@@
   • All requests served (11) ensures that all requests during each interval $\mathcal{T}_{i}$ are attributed to a quality tier.
-  • Sufficient resources (11) ensures that the provided machines have sufficient capacity to serve the requests allocated to each quality tier.
+  • Sufficient resources (12) ensures that the provided machines have sufficient capacity to serve all requests per quality tier.
+  • Respect SLOs (13) ensures that all SLOs are respected.
   • Sufficient QoR (14) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
-Note, that this model assumes perfect knowledge of the future (in particular perfect forecasts for all $C^{i}$ and $R^{i}$) and can therefore not be solved in practice.
-However, we use it as an upper bound for assessing potential and as the foundation of the online proposed in Section 4.
-Complexity
-From a cloud tenant perspective (Scope 3), the optimization problem is an NP-hard mixed-integer linear program (MILP), as the objective and all constraints are linear, but involve the discrete decision variables $D^{i}$.
-From the perspective of data center operators, the Scope 2 power model from Equation (5) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$.
+Note, that this model assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$) and can therefore not be solved in practice.
+However, we use it for assessing upper-bound potential and as the foundation of the online approach proposed in Section 4.
+3.6. Complexity
+The proposed optimization problem is an NP-hard mixed-integer linear program (MILP), as we minimize a linear objective function subject to linear constraints, which include discrete decision variables in $D^{i}$.
+Equivalence of power models.
+From the perspective of data center operators, the load-dependant power model from (6) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$.
 This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve.
-However, during optimization, we can relax Equation (5) to the same form as Equation (7), by assuming $p^{i}_{m,q}=p_{m,q}^{\text{max}}$, as long as $n\leq 1$, i.e.
+However, during optimization, we can relax (6) to the same form as (8), as long as $n\leq 1$, i.e.
 the power model is concave.
-If the power law relationship is strictly convex, i.e.
-$n>1$, it becomes more efficient to deploy multiple machines at lower utilization levels $util^{i}_{q}$ rather than fully utilizing fewer machines.
+Theorem 3.1.
+We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$.
+When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
+$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
+can be simplified to
+$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
+as it always yields an equivalent optimal deployment $D^{i}$.
+The proof is provided in Appendix A.
+Intuitively, only if the power-law relationship is strictly convex ($n>1$), it can be more efficient to deploy multiple machines at lower utilization rather than fully utilizing fewer machines.
 However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007).
-Hence, for $n\leq 1$, the power models for Scope 2 and Scope 3 reporting are equivalent during optimization, as they result in the same decisions for optimal $A^{i}$ and $D^{i}$.
-Lastly, when values in $D^{i}$ are sufficiently big (i.e.
-in large-scale scenarios with many active machines), the optimal result can be approximated by relaxing $D^{i}$ to continuous variables.
-This turns the problem into a simple linear program that is solvable in polynomial time.
+Hence, for $n\leq 1$, we can use a load-independent power model during optimization, as it yields the same optimal $D^{i}$.
 4. Online Approach
-For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$ as time progresses.
+For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$.
 Its performance is limited by two factors:
   1. (1)
-    the errors of request and carbon intensity forecasts
+    the error of request and carbon intensity forecasts
   2. (2)
-    the required approximations for solving the NP-hard optimization problem
-Yet, our experiments show that even under realistic forecasts and strict time limits for solving the MILP, we can achieve close-to-optimal emission savings (Section 5.4).
-4.1. Forecasting
-We denote forecasts for request volume and carbon intensity as $\widehat{R}^{i}$ and $\widehat{C}^{i}$, respectively.
-There are a variety of different approaches commonly for long-term forecasts with multiple seasonalities, like Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011).
-As the electricity mix of regions does usually not change abruptly, long-term forecasts for ACI can be estimated reasonably well by fitting daily, weekly, and yearly seasonalities.
-However, especially in regions that heavily rely on weather-dependent renewable power production, additionally utilizing short-term forecasts (typically up to 4 days) provides significantly better accuracy (Maji et al., 2023; Li et al., 2024c).
+    the error introduced by approximating the NP-hard optimization problem
+In our experiments, conducted with realistic forecasts and strict time limits for solving the MILP, we achieve emission savings of $82\pm 6$ relative to the upper-bound potential (Section 5.4).
+4.1. Multi-Horizon Optimization
+Based on the optimization problem introduced in Section 3.5, we propose a forecast-based multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
+Forecasting.
+There exist various approaches for long-term load forecasting, such as Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011).
+As the electricity mix of regions does not change abruptly, long-term forecasts for ACI can be usually estimated reasonably well by fitting daily, weekly, and annual seasonalities.
+However, especially in regions that are heavily dependent on weather-dependent renewable power production, additionally using short-term forecasts can provide significantly higher accuracy (Maji et al., 2023; Li et al., 2024c).
 For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
-4.2. Multi-Horizon Optimization
-Based on the optimization problem introduced in Section 3.5, we introduce a multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
+Algorithm.
 The approach is described in Algorithm 1.
+As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality.
+To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
 Algorithm 1 Multi-Horizon Optimization
@@
 6:    # Short-term optimization
-7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}}\sum_{i=\alpha}^{I-1}E^{i}$
+7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}}\sum_{i=\alpha}^{I-1}E^{i}$
 8:    # Progress in time
@@
 10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
-As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality.
-To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
   1. (1)
@@
 In Line 7 we assume $\gamma$ as the horizon..
-As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly.
+As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly.
 If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (12) constraint.
@@
 After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
-4.3. Dynamic QoR Adaptation for Best-Effort Users
+4.2. Automatic QoR Adaptation
 The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$.
-However, especially for best-effort users, another relevant question is: What are the expected yearly emissions from this user group, and can we keep them within a predefined budget?
-In Section 5.5, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected yearly emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year.
+However, especially for best-effort users, another relevant question is: What are the expected annual emissions from this user group, and can we keep them within a predefined budget?
+In Section 5.4, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected annual emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year.
 Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:
 (15) $\max\quad$ $\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
-[…]
 […]
@@
 (17) $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\max_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\left(\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=0}^{I-1}\right)$
-Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate the emissions from all intervals in Equation (14).
+Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate emissions over all intervals in (16).
 This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$.
@@
 5. Experiments
-We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the entire year 2023 using two real-life request traces and in ten different regions.
-All experiments were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job. We use Gurobi (Gurobi Optimization, LLC, 2024) for solving the optimization problems.
+We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
+We use Gurobi (Gurobi Optimization, LLC, 2024) to solve optimization problems. All simulations were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job.
+Table 2.
+Upper-bound potential and carbon savings under realistic conditions in % for all request traces and regions at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$.
+Note, that other $\text{QoR}_{\text{target}}$ can lead to even better results (Figure X).
+Region Artificial Real Synthetic Mean
+Static Random Wiki (en) Wiki (de) Taxi Cell B Cell D Cell F
+NL 8.2 / 7.4 8.5 / 6.3 7.6 / 6.5 10.2 / 9.3 11.3 / 8.7 7.8 / 6.7 7.6 / 6.6 8.0 / 7.0 8.7±1.3 / 7.3±1.0
+CISO 7.0 / 6.8 7.5 / 6.0 7.5 / 7.0 10.0 / 9.4 11.3 / 9.2 6.6 / 6.1 6.3 / 5.9 6.8 / 6.4 7.9±1.7 / 7.1±1.3
+ES 7.1 / 6.3 7.3 / 5.2 6.5 / 5.6 9.4 / 8.4 10.4 / 8.0 6.6 / 5.6 6.6 / 5.7 6.9 / 5.9 7.6±1.4 / 6.3±1.1
+AU-QLD 7.9 / 7.8 8.4 / 7.0 7.1 / 6.5 6.1 / 5.8 8.8 / 6.9 7.6 / 7.0 6.9 / 6.4 8.0 / 7.5 7.6±0.8 / 6.9±0.6
+DE 6.3 / 5.4 6.7 / 4.5 5.7 / 4.6 8.7 / 7.6 9.7 / 6.9 5.9 / 4.8 5.8 / 4.8 6.1 / 5.1 6.9±1.4 / 5.5±1.1
+PL 4.1 / 3.8 4.5 / 2.9 3.5 / 2.9 6.7 / 6.2 7.5 / 5.7 3.8 / 3.3 3.7 / 3.2 4.1 / 3.5 4.7±1.4 / 3.9±1.2
+ERCOT 4.2 / 3.1 4.6 / 2.6 3.6 / 2.4 6.3 / 5.1 7.4 / 5.8 3.9 / 2.8 3.7 / 2.5 4.1 / 3.0 4.7±1.3 / 3.4±1.2
+SE 2.7 / 2.4 3.0 / 1.6 2.0 / 1.5 4.9 / 4.5 6.1 / 5.4 2.3 / 1.9 2.2 / 1.7 2.6 / 2.2 3.2±1.4 / 2.6±1.4
+NYISO 2.3 / 1.8 2.7 / 1.1 1.7 / 1.1 4.6 / 4.3 5.6 / 4.6 1.9 / 1.5 1.8 / 1.3 2.2 / 1.7 2.8±1.4 / 2.2±1.3
+PJM 1.8 / 1.7 2.3 / 0.8 1.3 / 1.0 4.1 / 4.0 5.5 / 4.3 1.6 / 1.4 1.4 / 1.1 1.8 / 1.6 2.5±1.4 / 2.0±1.3
 5.1. Experimental Setup
-As average carbon intensity is usually reported hourly, we consider $\Delta=1$ hour.
-Scenario
+As ACI is usually reported hourly, we define $\Delta=1$ hour.
 […]
 […]
-We model machine performance using recent benchmarks of vLLM on a EC2 p4d.24xlarge for inference, which state a throughput of 11.57 requests per second for LLaMA 3.1 8B and 5.05 requests per second for LLaMA 3.1 70B (vLLM Team, [n.
-d.]).
-Request traces
-We simulate realistic request patterns using two hourly request traces from the Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), as shown in Figure 5. In particular, we use
-  • Wiki (en): the English Wikipedia with $63.4e6\pm 0.8e$ 6 requests per hour and spikes of up to $616.4e$ 6 requests per hour.
-In our scenario, this results in 80-200 active machines throughout the year.
-  • Wiki (de): the German Wikipedia with $54.7e5\pm 2.8e$ 5 requests per hour, with spikes of up to $517.6e$ 5 requests per hour.
-This results in only 10-25 active machines, emphasizing the importance of discrete scheduling decisions.
-Figure 5. The request traces differ notably, as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
-Carbon intensity regions
-We analyze the potential in ten different regions:
+Request traces.
+For our analysis, we require 4 years of hourly request traces: 2020-2022 for fitting forecasting models and 2023 for the analysis itself.
+As such data are not publicly available for LLM inference servies, we show the effectiveness of our method across a diverse set of other real and synthetic request traces:
+  • Artificial: A Static trace that assumes a constant stream of hourly requests, and a Random trace, where we sample hourly requests from a normal distribution.
+  • Real: Two real-life traces representing the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n.
+d.]), Wiki (en) and Wiki (de), and a trace based on the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
+  • Synthetic: Three generated traces, called Cell B, D, and F, which are based on instance event traces of Borg cluster cells at Google (Google, 2019).
+All request trace datasets and our forecast methodology are explained in detail in Appendix B.
+Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
   • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
   • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
-  • In Australia, we consider Queensland (AUS-QLD)
+  • In Australia, we consider Queensland (AU-QLD)
 All carbon intensity data was provided by ElectricityMaps (Maps, 2024).
-We normalize all time zones to align patterns with the time of day in the request traces.
-Temporal variability of hourly carbon intensities are shown in Figure 1.
-5.2. Impact of QoR on Energy Demand
-We first analyze the absolute yearly emissions generated by our different scenarios without carbon-aware QoR adaption.
-For this, we solve Equation (9) for both request traces, across all 10 regions, $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$, and $\gamma=1$ hour.
-The results reflect emission savings driven purely by energy demand, i.e., by serving more or less requests on Llama 3.1 70B.
-We optimize until the objective is within 0.1 % of its optimal solution or after 10 minutes.
-All results are within 0.8 % of the optimal solution.
-Figure 6 shows the overall emissions for Wiki (en).
-As expected, we observe very large differences across different regions.
-For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally.
-Serving all requests with the 80B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source.
+We normalized all time zones to align the time of day with the request traces.
+The temporal variability of regions is depicted in Figure 1.
+Our forecast methodology is explained in Appendix C.
+5.2. Absolute Emissions
+We first analyze the absolute annual emissions without carbon-aware QoR adaption, i.e.
+for $\gamma=1$ hour.
+The results reflect emission savings driven purely by increased energy demand from serving more requests on Llama 3.1 70B ($\text{QoR}_{\text{target}}\to 1$).
+Figure 4 shows the annual emissions for Wiki (en)—the trend is equivalent for all datasets.
+We observe very large differences across different regions: For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally.
+Serving all requests with the 70B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source.
 As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$.
-Results are equivalent for Wiki (de), but smaller by a factor of around 7.5.
-Figure 6. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ for all 10 regions.
-Figure 7.
-Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$.
-Bigger validity periods yield larger gains due to increased flexibility.
-5.3. Upper-Bound Potential of Carbon-Aware QoR Adaptation
+Figure 4. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
+Figure 5. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains.
+5.3. Upper-Bound Potential
 […]
-Carbon Savings Beyond Energy Efficiency
-As define scenarios with validity periods of $\gamma=1$ hour as baselines.
-For these scenarios, emissions are directly correlated with energy demand as there is no flexibility for carbon-aware QoR adaptation.
-Figure 7 reports the relative and absolute additional savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods with respect to these baselines.
+Carbon savings beyond energy efficiency.
+In blue (left), Table 2 presents the relative additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week compared to $\gamma=1$ hour, where emissions directly depend on energy demand.
+We observe that through carbon-aware QoR adaptation, emissions drop by around 8 % in many regions.
+While request trace patterns exhibit little difference in potential, traces with a low volume of requests (Wiki (de) and Taxi) show increased potential, due to the higher impact of discrete scheduling decisions.
+For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines.
+In experiments with even fewer machines we managed to gain relative savings of more than 40 %—however, we do not analyze such cases further, as we focus on high-impact scenarios with many machines.
+Impact of validity periods.
+Figure 5 reports the relative and absolute savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods.
 Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings.
 This is because average carbon intensity shows limited short-term variation, leaving little room for optimization.
-However, starting at $\gamma=24$ hours, we can see more notable potential of 5-8 % in some regions.
-The characteristics of carbon intensity data significantly influence the expected gains.
-For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany experiences more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021).
-As a result, extending $\gamma$ beyond 24 hours has minimal impact in California, whereas in Germany, larger $\gamma$ values lead to considerable improvements.
-In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity, which limits the potential for carbon-aware approaches.
-Impact of $\text{QoR}_{\text{target}}$
-Figure 8. Additional relative savings for different $\text{QoR}_{\text{target}}$ in Germany and California.
+However, starting at $\gamma=24$ hours, we see a more notable potential of 5-8 % in some regions.
+Regional characteristics significantly influence the expected gains: For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany exhibits more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021).
+As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements.
+In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity.
+Impact of premium users.
+Figure 6 illustrates the loss of potential with a growing fraction of premium users at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$.
+We observe a slightly sub-linear decrease with a certain offset which depends on the overall request volume and, hence, the impact of discrete scheduling decisions.
+For reference, ChatGPT reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024).
+The resulting 5 % premium users decrease the potential savings by 5.6 % on average.
+Figure 6. Impact of premium users on the potential savings.
+Impact of $\text{QoR}_{\text{target}}$.
 Figure 7 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$.
-When $\text{QoR}_{\text{target}}$ is 0, the optimizer will serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings.
+At $\text{QoR}_{\text{target}}=0$, we serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings.
 Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B.
-Values around $\text{QoR}_{\text{target}}=0.5$ offer the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
-Figure 9. Additional carbon savings under realistic conditions. Our online approach maintains additional savings within one percentage point across most experiments.
-Impact of validity periods
-(a) $\beta=\{1,4,8,24\}$ hours for smaller validity periods.
-(b) $\beta=\{1,3,7\}$ days for larger validity periods.
-Figure 10.
-Considering all time intervals of length $\beta$ throughout the year, what proportion of these intervals have a QoR below the threshold?
-We analyze the CDF for Wiki (en) in Germany for $\text{QoR}_{\text{target}}=0.5$.
-While increasing $\gamma$ enhances potential for carbon savings, it also raises the risk of prolonged periods with low QoR.
-Service providers can adjust the range of QoR using SLO and SLG to ensure that a certain fraction of requests can always be served from a specific quality tier.
-However, limiting the fraction of requests available for QoR adaptation lowers its potential impact.
-Figure 10 illustrates the probability of prolonged periods with low QoR under different validity period lengths.
-To to provide an example of how to interpret the results: Figure 10(a) shows that when optimizing for a $\text{QoR}_{\text{target}}=0.5$ with validity periods of $\gamma=8$ hours, no 8-hour interval ($\beta$) has a QoR below 0.5, while all intervals maintain a QoR of at least 0.5.
-However, around 10 % of all 4-hour intervals fall below a QoR of 0.25.
-We observe, that for validity periods of $\gamma=3$ months, users can experience several consecutive days of low QoR.
-Consequently, we conclude that very large validity periods of more than a month are not practical in real-world scenarios.
-5.4. Potential under Realistic Conditions
-We evaluate the practicability of carbon-aware QoR adaptation using our proposed online optimization under real forecasts and near-optimal MILP solutions.
-Setup
-We simulate request forecasts using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and yearly seasonalities.
-In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year.
-The predictions exhibit realistic errors: For Wiki (en), 24-hour forecasts have an average MAPE of $13.9\pm 8.4$, while for Wiki (de), the error is $32.1\pm 15.3$.
-For long-term carbon intensity forecasts, we predict the entire year of 2023 as of January 1st using Prophet, fitted on three years of historical data.
-We simulate short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region (see Appendix A).
-Short-term forecasts are updated daily at midnight.
-We interrupt long-term optimizations after 30s and short-term optimizations after 10s, to demonstrate performance under near-optimal solutions.
-Furthermore, we limit this analysis to daily, weekly, and monthly validity periods, as Section 5.3 showed that validity periods shorter than a day offer minimal potential for carbon savings, while periods longer than a month can lead to extended durations of low QoR.
-Results
-Figure 9 shows the performance gap of our online optimization compared to the theoretical upper bound established in Section 5.3.
-Across all load traces and regions, the resulting emissions savings are generally within one percentage point of the optimal solution.
-In regions with already limited upper-bound potential, the expected savings become more marginal.
-However, in most regions, our approach consistently maintains additional savings of 5–10 %.
-Within the stated time limits, 95 % of all long-term optimizations achieve an optimization gap $<2.1\,\%$, while 95 % of all short-term optimizations reach a gap $<0.3\,\%$.
-5.5. Outlook of Yearly Emissions
-(a) Yearly emission outlook at three discrete points in time.
-(b) Yearly emission outlook, performed weekly for $\text{QoR}_{\text{target}}=0.5$.
-Figure 11.
-Projected yearly emissions (blue) based on long-term forecasts for different $\text{QoR}_{\text{target}}$.
-The interval shows the range between best- and worst-case forecasts of $\widehat{R}$, while red indicates emissions already incurred.
-Long-term optimizations become computationally cheap when accepting near-optimal solutions.
-This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts $\widehat{R}$, which may represent, for instance, worst-case or best-case scenarios.
-Setup
-We demonstrate emission outlooks by executing the online approach on the Wiki (en) traces in California.
+A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
+Figure 7. Additional relative savings for different $\text{QoR}_{\text{target}}$.
+Figure 8. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?
+Periods of low service quality.
+While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 8.
+For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right).
+However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left).
+We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR.
+Consequently, very large validity periods of more than a month are not practical in real-world scenarios.
+5.4. Evaluation under Realistic Conditions
+We evaluated the practicability of carbon-aware QoR adaptation using the proposed multi-horizon optimization under real forecasts (see Appendix B as well as C) and approximated MILP solutions: Long-term optimizations are stopped at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 30s; short-term optimizations at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 10s.
+[…]
+Figure 9. Additional carbon savings under realistic conditions for Wiki (de) across all regions and $\text{QoR}_{\text{target}}=0.5$.
+Outlook of annual emissions.
+Long-term optimizations are computationally cheap when accepting near-optimal solutions—within just 30 seconds, the mean $\text{MILP}_{\text{gap}}$ reached 0.53 % accross all experiments.
+This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts, to represent, for instance, worst-case or best-case scenarios.
+Figure 10 demonstrates emission outlooks for Wiki (en) traces in California.
 For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively.
 Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
-Results
+(a) Annual emission outlook at three discrete points in time.
+(b) Annual emission outlook; weekly for $\text{QoR}_{\text{target}}=0.5$.
+Figure 10. Projected emissions (blue) based on long-term forecasts. The interval shows the range between best- and worst-case forecasts. Red indicates emissions already incurred.
 Figure 10(a) shows the resulting outlooks a the start of the year, after 10 weeks, and after 40 weeks.
@@
 Figure 10(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$.
-We observe that the upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as actual emissions accumulate.
-5.6. Dynamic QoR Adaptation for Best-Effort Users
-Figure 12. Daily QoR for California, Poland, and Sweden. Our adaptive approach provides the most consistent QoR throughout the year.
-Setup
-We illustrate the dynamic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.3 on Wiki (en) in California, Poland, and Sweden to cover a spectrum of differently behaving carbon intensity traces: California exhibits high variability due to many solar deployments, Poland has relatively little variation due to reliance on fossil fuels, and Sweden usually has very low carbon intensity which occasionally spikes when demand needs to be met through cross-border imports or natural gas.
-For each region, we define the annual budget $B^{\text{annual}}$ so that the ideal $\text{QoR}_{\text{target}}=0.50$.
-We compare our approach to two greedy baselines:
+The upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as incurred emissions accumulate.
+5.5. Automatic QoR Adaptation
+Figure 11. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL). Our approach provides the most consistent QoR throughout the year.
+Lastly, we illustrate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.2 on Wiki (en) in three exemplary regions.
+For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment (Section 5.3) for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$.
+Consequently, under perfect forecasts, the optimal approach would be to choose $\text{QoR}_{\text{target}}=0.5$ for every time step.
+We compare our approach to two greedy baselines that have access to the same forecasts as our approach:
   1. (1)
-    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves Equation (13).
-Any remaining or overspent budget is evenly distributed across the remaining time steps.
+    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves (15). Any remaining or overspent budget is evenly distributed across the remaining time intervals.
   2. (2)
-    Greedy (weighted) solves Equation (13), too, but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as forecasts are updated, weightings get adjusted.
-Results
-Figure 12 illustrates the daily QoR of our online approach compared to both greedy baselines over time.
-Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR fluctuates significantly throughout the year.
-Greedy (weighted) delivers more consistent QoR for both Poland and Sweden, though it starts very conservatively in both regions, leading to near-perfect QoR at the year’s end due to unused budget.
-In California, where forecasts are less accurate, this baseline overspends early in the year, resulting in days with $\text{QoR}=0$ for most of the year’s end.
-Our approach also starts conservatively, as the forecast initially overestimates annual emissions (see Figure 11).
-However, it delivers the most consistent service across all experiments.
+    Greedy (weighted) also solves (15), but weights the hourly budgets based on forecasted load and carbon intensity.
+Every 24 hours, as request and carbon intensity forecasts are updated, weightings get adjusted.
+Figure 11 shows the daily QoR of the greedy baselines and our automatic QoR adaptation.
+Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year.
+Greedy (weighted) delivers more consistent QoR in all experiments but tends to fall back to either $\text{QoR}=1$ (NL) or $\text{QoR}=0$ (CISO and NYISO) toward the end of the year due to forecast errors.
+In the CISO and NYISO regions, it even exceeds the budget by 0.8 and 1.2 %, respectively.
+Our approach provides the most consistent service across all experiments as it periodically adjusts $\text{QoR}_{\text{target}}$ based on long-term optimizations.
+The annual budget is utilized to 98.2–99.8 %.
 6. Discussion
-We highlight the future impact of carbon efficiency in energy-intensive services and line out the limitations of this study.
-Impact
-We want to emphasize the extend of carbon efficiency improvements for cloud services, based on our analysis in Section 4.3.
-ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024).
-Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our scenarios.
+This section highlights the impact of our results on energy-intensive services and discusses the limitations of this study.
+Impact.
+We emphasize the significant potential for carbon efficiency improvements in cloud services, based on our analysis in Section 5.2.
+ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024).
+Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our larger scenarios.
 Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
 Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy.
-However, and increasing part of the community is advocating to base load shifting decisions based on marginal data (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids.
-However, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone.
+However, and increasing part of the community is advocating to base load shifting decisions based on marginal signals (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids.
+Yet, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone.
 Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion.
-We predict, that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods.
-Limitations
-In this paper, we provide first insights into the potential of carbon-aware QoR adaptation based on an LLM inference use case.
-Applying our approach to more complex, real-world scenarios with multiple user groups, service quality tiers, and heterogeneous infrastructure is possible from a modeling perspective, but not yet investigated.
-Furthermore, we only concentrate on QoR within a single data center and do not distribute load across multiple geo-distributed regions—which is known to yield the largest potential for carbon savings (Sukprasert et al., 2024b).
-This paper represents the first comprehensive analysis of QoR adaptation for energy and carbon efficiency, and sets the stage for more detailed explorations in future research.
+We predict that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods.
+Our approach is directly applicable to MCI signals or any other scalar metric without requiring adaptations.
+Limitations and future work.
+This paper presents the first analysis on the potential of carbon-aware QoR adaptation in services, focusing on an LLM inference use case.
+While our approach is already applicable to more complex scenarios with multiple service quality tiers and heterogeneous infrastructure, such cases have not yet been investigated experimentally.
 Second, our current modeling only considers the QoR for user groups, rather than individual users.
@@
 For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group.
-We consider addressing fairness aspects with more granular user modeling as as an area for future work.
+We consider fairness aspects with more granular user modeling as an area for future work.
+Third, future work should explore the impact of adaptive QoR adjustments on user quality of experience and the resulting implications on user behavior.
+For instance, in the context of generative AI, reduced QoR might incentivize users to resubmit requests, potentially offsetting carbon savings.
+Unfortunately, to the best of our knowledge, there are no established behavioral models for generative AI usage.
+Developing such models would be valuable to better quantify the real-life impacts of sustainability approaches, including rebound effects.
 7. Related Work
 We survey related works on carbon-aware computing as well as carbon efficiency in energy-intensive cloud services.
-Carbon-aware computing
+Carbon-aware cloud services.
 The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy.
-Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024; Sukprasert et al., 2024b).
+Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a; Lechowicz et al., 2024).
 However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
 Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed data centers and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016).
-For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint while respecting latency constraints.
+For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint under latency constraints.
 CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks.
 Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity.
-However, no existing approaches address services that must operate in specific data centers due to data, security, or infrastructure constraints.
-Sustainability of cloud services
-Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023).
-Especially the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024).
-While service degradation is a common practice to prevent overload (Meza et al., 2023; Park et al., 2024), it has already been proposed to design applications to offer different quality tiers of a service for sustainability (Vitali, 2022).
-In the context of carbon-aware LLMs, there only exists a single approach, Sprout (Li et al., 2024b), which optimizes an LMMs autoregressive generation process.
-In particular, the authors propose to return shorter responses during periods of high carbon intensity, as the energy demand of a response directly correlates with the amount of generated tokens.
-Sprout clearly classifies as an approach which adapts QoR over time and can benefit from our scheduling decisions.
+However, there are no existing approaches that address services which must operate in a single region due to data, security, or infrastructure constraints.
+Sustainability in generative AI inference.
+Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023; Samsi et al., 2023).
+Especially, the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024).
+Many recent approaches target sustainability aspects of inference services throuh power-aware scheduling (Stojkovic et al., 2024, 2025; Qiu et al., 2025; Patel et al., 2024) or by considering their carbon and water footprint in geographic load balancing (Li et al., 2024d).
+Service quality adaptation.
+While service degradation is a common strategy to prevent system overload (Meza et al., 2023; Park et al., 2024), recent research has proposed designing applications with different service quality tiers as a means of improving sustainability (Vitali, 2022).
+In the context of carbon-aware LLMs, we identified only one approach, called Sprout (Li et al., 2024b), which optimizes the autoregressive generation process.
+Sprout returns shorter responses during periods of high carbon intensity, as the energy demand directly correlates with the number of generated tokens.
+Thereby, it effectively adapts QoR over time and can be further enhanced by leveraging the optimized infrastructure deployments and load balancing proposed in this paper.
 8. Conclusion
@@
 In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency.
-Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand alone.
-Moreover, we propose an approach for dynamically adjusting the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
+Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand.
+Moreover, we propose an approach to automatically adapt the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
 Acknowledgements.
 This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. 01IS18025A and 01IS18037A).
-Appendix A CarbonCast MAPE
+Appendix A Proof of Theorem 3.1
+Theorem 3.1 We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$.
+When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
+$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
+can be simplified to
+$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
+as it always yields an equivalent optimal deployment $D^{i}$.
+Lemma A.1.
+Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines than strictly necessary.
+Formally, if $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$, an utilization-dependent power model $f$ is strictly subadditive
+(18) $f(x+y)<f(x)+f(y),$
+where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
+Proof of Lemma A.1.
+We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
+Case 1: Linear power model ($n=1$):
+(19) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
+$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
+(20) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
+(21) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
+(22) $f(x+y)$ $<f(x)+f(y)$
+Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $f(0)\geq 0$, $f$ is strictly subadditive on $(0,\infty)$.
+(23) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
+(24) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
+(25) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl{(}}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl{(}}}_{>1}<0$
+(26) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
+Additionally, $f(0)=p_{m,q}^{\text{max}}>0$. ∎
+Proof of Theorem 3.1.
+For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
+[…]
+  1. (1)
+    Feasibility with fewer machines is impossible. If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate Sufficient resources (12).
+  2. (2)
+    Using extra machines is suboptimal.
+By Lemma A.1, under the utilization-dependent power model, there is no incentive to deploy extra machines.
+Likewise, any extra machine increases the objective under the simplified power model.
+We conclude that, for both power models, any optimal solution must yield an equivalent deployment $\check{D}^{i}$ for each interval $i$. ∎
+Appendix B Request Traces
+This section describes all request trace datasets used in the experiments. The traces are visualized in Figure 12; Table 3 presents summary statistics.
+Table 3. Request trace and 24-hour forecast statistics.
+Dataset statistics ($1\times 10^{6}$) Forecast statistics
+Mean ± Std Dev Min Max MAPE
+Static 1.00 ± 0.00 1.00 1.00 0.0 ± 0.0
+Normal 1.00 ± 0.34 0.00 2.36 38.6 ± 24.6
+Wiki (en) 3.38 ± 0.80 1.88 16.41 13.9 ± 8.4
+Wiki (de) 0.42 ± 0.24 0.04 1.56 32.1 ± 15.3
+Taxi 0.33 ± 0.14 0.04 0.71 26.5 ± 7.1
+Cell B 1.94 ± 0.61 0.73 4.10 27.2 ± 13.5
+Cell D 2.87 ± 0.80 1.02 7.76 22.1 ± 15.8
+Cell F 1.58 ± 0.41 0.87 4.32 18.2 ± 9.3
+Figure 12. First month of data and KDE over all data points in 2023 for the eight datasets used in the evaluation.
+Artificial Datasets.  We artificially create two traces:
+  • Static assumes a constant stream of $1\times 10^{6}$ hourly requests.
+  • Random comprises points randomly sampled from a normal distribution with a mean of $\gamma=1\times 10^{6}$ and a standard deviation of $\sigma=0.33\times 10^{6}$.”
+During online optimization, we always assume the mean of $1\times 10^{6}$ requests per hour as forecasts.
+Real Datasets.  We create three traces based on real data:
+  • Wiki (en) is based on the Wikipedia pageview statistics (Wikimedia Foundation, [n. d.]), namely all hourly requests to en.wikipedia.org.
+  • Wiki (de) represents all requests to de.wikipedia.org.
+The trace differs notably from Wiki (en), as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
+  • Taxi is based on hourly aggregates of taxi trips in New York City (Taxi and Commission, 2025).
+We sum over all trip records (Yellow Taxi, Green Taxi, For-Hire Vehicle, High Volume For-Hire Vehicle) and multiply the events by factor 10, to bring them into a similar range with the other datasets.
+To simulate realistic forecasts, we are using using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and annual seasonalities.
+In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year.
+The predictions exhibit realistic errors, denoted in Table 3.
+Synthetic Datasets.
+As all real datasets exhibit very periodic patterns, we additionally created three synthetic datasets to demonstrate that carbon-aware QoR adaptation is feasible on highly unpredictable traces.
+For this, we aggregated the instance events of the Google cluster traces (version 3) (Google, 2019) individually for the eight available Borg cells.
+Data was retrieved via the GoogleSQL query:
+SELECT
+  TIMESTAMP_TRUNC(TIMESTAMP_MICROS(time), HOUR) AS hour,
+  COUNT(*) AS num_events
+FROM ‘google.com:google-cluster-data.<TABLE>‘
+-- Exclude invalid values
+WHERE time > 0 AND time < 9223372036854775807
+GROUP BY hour
+ORDER BY hour;
+[…]
+We identified the three traces with the lowest 24-hour autocorrelation coefficients, i.e.
+the traces that exhibit the least daily seasonality: Cell B (0.17), Cell D (0.27), and Cell F (0.22).
+Since the Google cluster dataset only covers one month of data, we fit a DoppelGANger (Lin et al., 2019) model for each of the traces, which is a state-of-the-art GAN for timeseries generation, and generate 4 years of synthetic data per cell.
+We apply the same forecasting methodology that is used for real datasets.
+Appendix C Carbon Intensity Forecasts
+For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets.
+As in (Lechowicz et al., 2024), we generate synthetic short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region, see Table 4.
+Short-term forecasts are updated daily at midnight.
+Table 4. Carbon forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
 Region Day 1 Day 2 Day 3 Day 4
@@
 AU-QLD 3.93 3.98 4.06 5.87
-Table 2. Carbon forecast MAPE over 96 hours for all observed regions as reported in CarbonCast (Maji et al., 2023)

### v3

@@
-Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services
+Carbon-Aware Quality Adaptation for Energy-Intensive Services
 Abstract.
 The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace.
-While hyperscalers collectively fail to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions.
 To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing.
@@
 In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers.
-We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency.
-Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget.
-Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
+We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency and introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings.
 Sustainable computing, quality of service, LLM inference, green AI
-submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v2. Under review.
+journalyear: 2025copyright: ccconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computing
 1. Introduction
 Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023).
-Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024).
+Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024a; Wu et al., 2024).
 As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
-Figure 1. The carbon intensity of electricity can vary significantly over time. Exploiting these temporal variations can reduce the carbon footprint of energy-intensive services.
-Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi).
-While reporting on Scope 2 emissions (GHG emissions caused by electricity consumption) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI), upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023)—including emissions from the use of cloud services.
-This will affect not only service providers, but also users of energy-intensive services.
 The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is.
-Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a).
-However, for continuously running services, where individual requests have little to no delay-tolerance, existing approaches exclusively focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013).
+Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024; Hanafy et al., 2024a).
+In contrast, existing carbon-aware solutions for interactive services, which handle individual requests with minimal delay-tolerance, predominantly focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013).
 However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations.
-For instance, generative AI inference typically relies on high-performance GPUs or specialized accelerators, which are often only available at specific locations.
 Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
-Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
-In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings.
-We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by different service quality tiers.
-For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021).
-Our approach minimizes the operational carbon emissions of energy-intensive services by optimizing resource provisioning and load balancing across service quality tiers, while ensuring compliance with service level objectives (SLOs) for different user groups.
+Figure 1. We investigate how adjusting the quality of responses (QoR) with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
+In this work, we focus on the temporal variability of carbon intensity and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings.
+We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by two different service quality tiers.
+For instance, the service quality and carbon footprint of LLM inference is determined by many factors such as the choice of model (Faiz et al., 2024), quantization and pruning (Liang et al., 2021), and the number of generated tokens (Li et al., 2024a).
 Figure 1 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week.
-Serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0).
+Here, serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0).
 Balancing requests evenly between the two tiers results in intermediate energy usage.
 We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints.
-Based on this, we propose an approach that achieves close-to-optimal savings despite real-life challenges such as forecast inaccuracies and the high computational complexity of the underlying optimization.
 Contributions.  Towards more carbon-efficient cloud services, we make the following contributions:
   1. (1)
-    We formalize the problem of minimizing carbon emissions under QoR constraints through optimized resource provisioning and load balancing between service quality tiers.
+    We formalize the problem of minimizing emissions under QoR constraints through optimized resource provisioning and load balancing between two service quality tiers.
   2. (2)
-    We model emissions from the perspective of data center operators and cloud tenants (Scope 2 and 3 reporting) and prove that both models lead to equivalent decisions.
+    We propose a forecast-based multi-horizon optimization for minimizing emissions that reaches close-to-optimal carbon savings.
   3. (3)
-    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can automatically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
-  4. (4)
-    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service on eight real and synthetic request traces in ten different regions.
-All experiments in this paper are fully reproducible. All code and data will be made publicly available upon acceptance.
-2. Background and Motivation
-Despite previous commitments to achieve fully carbon-free or even “carbon-negative” operations by 2030, Microsoft recently announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), Google reported a 48 % increase in GHG emissions within five years (Google, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI.
-The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
-All big hyperscalers have announced or introduced carbon reporting tools for their customers (Google Cloud, 2023; Microsoft, 2023; Amazon Web Services, 2023; Dixit and Tse, 2024).
-However, they are expected to significantly underreport emissions, due to opaque and inconsistent reporting methodologies (O’Brien, 2024).
-We anticipate an increasing focus on regulatory frameworks that will require cloud providers to disclose their carbon emissions more comprehensively.
-Energy-intensive cloud services.
-The rapid adoption of large language models (LLMs) requires immense processing power, not only for training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024).
-As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase future electricity demand (Li et al., 2024a).
-Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands.
-Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024).
-Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024).
-Additionally, scientific simulations such as climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022) are growing in complexity and increasingly integrate AI services.
-Growing pressure from GHG reporting.
-The reporting of Scope 2 emissions, i.e.
-indirect GHG emissions caused by consuming electricity, is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
-In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024).
-While, the disclosure of Scope 3 emissions—i.e.
-all indirect emissions that occur across the value chain—stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023).
-Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019).
-From 2026 on, this will also include non-EU companies with significant operations in the EU.
-As the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018), also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
-Implications for energy-intensive services.  When reporting emissions from the perspective of cloud service providers, we must distinguish between two cases:
-  1. (1)
-    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity usage, i.e. Scope 2 emissions.
-  2. (2)
-    They rent infrastructure from cloud providers, in which case the data center’s emissions will be passed down the value chain als Scope 3 emissions, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
-Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023).
-We will address the two cases in our problem formalization, however, both are affected by stricter GHG reporting.
-Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost.
-Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
-Assumptions on future regulatory policies.
-As the effectiveness of annual and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions.
-Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
-We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024b; Radovanovic et al., 2022).
-ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source.
-Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time.
-Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework.
-In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
+    We evaluate the potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM service.
+All code and data is publicly available [^1] https://github.com/dos-group/qualitytime.
 2. Carbon-Aware QoR Adaptation
@@
 Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023).
-Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency.
-For instance, service quality tiers for LLMs can vary based on factors such as the choice of model (e.g., number of parameters), quantization (ranging from 4-bit integers to 32-bit floating-point), and the number of tokens generated (i.e.
-providing shorter replies (Li et al., 2024b)), and can have a significant impact on the carbon footprint of inference.
-In this work, we examine the impact of dynamically adjusting service QoR over time to enhance carbon efficiency within a single region, without relying on geo-distributed load balancing.
-3.1. Problem Setting
-We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time intervals
-(1) $\mathcal{T}_{i}=[i\Delta,(i+1)\Delta],\quad i=0,1,\dots,I-1,$
-where $I=T/\Delta$ is the total number of intervals, i.e.
-$i=0$ corresponds to the first interval and $i=I-1$ to the last.
-For the remainder of this paper, we consider $T=1~{}\text{year}$ and $\Delta=1$ hour.
-We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$ and service quality tiers $\mathcal{Q}$.
-Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
-For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group.
-Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$.
-The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$) [^1] Note, that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines.
-Furthermore, when elements in $D^{i}$ are sufficiently large (i.e.
-scenarios with many active machines), an optimal result can be approximated by relaxing $D^{i}$ to continuous variables.
-This turns the optimization problem introduced later into a linear program that can be solved in polynomial time..
-The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval.
-Table 1 shows an overview of all input and decision variables.
-Table 1. Description of symbols. Small letters denote the elements of the corresponding, capitalized matrix.
-Input variables
-$\mathcal{U}$ Set of user groups $u$
-$\mathcal{M}$ Set of machine types $m$
-$\mathcal{Q}$ Set of service quality tiers $q$
-$\gamma$ Validity period length
-$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $\mathcal{T}_{i}$.
-$C^{i}\in\mathbb{R}$ Carbon intensity during $\mathcal{T}_{i}$.
-$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
-$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
-$C_{m}^{\text{emb}}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
-Decision variables
-$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
-$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
-3.2. Quality of Responses
-We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers.
-In our subsequent analysis, we concentrate on a single user group, such as best-effort users, whose requests can be served by two distinct service tiers that differ in both response quality and energy demand.
-In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2.
+In this paper, we follow the predominant assumption that future regulations for Scope 2 emission reporting will be based on average carbon intensity (ACI), reported in carbon-equivalent emissions per unit of energy (gCO2/kWh) (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024; Radovanovic et al., 2022).
+We examine the impact of dynamically adjusting service QoR with respect to temporal changes in ACI to enhance carbon efficiency without relying on geo-distributed load balancing.
+[…]
+Quality of Responses.
+We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by two service quality tiers.
+That is, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2.
 When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
-In the following, we introduce several metrics to establish a more general definition of QoR, that holds for multiple user groups and service quality tiers.
-Service level objectives.
-Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), users of this framework can define a lower-bound service level objective $\underline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$, where $\underline{\text{\text{slo}}}_{u,q}$ indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$.
-Similarly, users can define an upper-bound $\overline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality.
-For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
-[…]
-$\underline{\text{SLO}}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{and}\quad\overline{\text{SLO}}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
-describes that premium users’ requests must always be served by Tier 2, while 60–100 % of requests by best-effort users can be served by Tier 1 and 0–40 % by Tier 2.
-All requests per user group must be served by a quality tier:
-(2) $\sum_{q\in\mathcal{Q}}\underline{\text{slo}}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\overline{\text{slo}}_{u,q}=1\quad\forall u\in\mathcal{U}.$
-Service level indicators.
-The service level indicator $\text{SLI}(\alpha,\omega)$ describes the observed distribution of requests across different quality tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
-(3) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
-As the service level is constrained by the SLOs, each
-(4) $\text{sli}_{u,q}(\alpha,\omega)\in\left[\min(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q}),\max(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q})\right].$
-General definition of QoR.
-We define the service’s quality of responses $\text{QoR}(\alpha,\omega)$ within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between $\text{SLI}(\alpha,\omega)$, $\underline{\text{SLO}}$ and $\overline{\text{SLO}}$.
-In particular, for the set of parameters that permit optimization, defined as $\mathcal{V}=\{(u,q)\in\mathcal{U}\times\mathcal{Q}:\overline{\text{slo}}_{u,q}\neq\underline{\text{slo}}_{u,q}\}$, we define QoR as:
-(5) $\text{QoR}(\alpha,\omega)=1-\min_{(u,q)\in\mathcal{V}}\left(\frac{\big{|}\overline{\text{slo}}_{u,q}-\text{sli}_{u,q}(\alpha,\omega)\big{|}}{\big{|}\overline{\text{slo}}_{u,q}-\underline{\text{slo}}_{u,q}\big{|}}\right).$
-Following the above example, an $\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$ would result in a $\text{QoR}(\alpha,\omega)=0.75$.
+Formally, for any interval $\bigcup_{i=\alpha,\dotsc,\omega}i$, QoR is defined as
+(1) $\text{QoR}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
 Validity periods.
-As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$.
-These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping.
+As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods of fixed length.
+These periods can either be disjoint (e.g., quarters, months, weeks) or, more commonly, overlapping.
 For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$).
@@
 Figure 2. Increasing validity periods allow for more flexibility in adjusting the proportion of requests served by Tier 1 and Tier 2 based on carbon intensity.
-3.3. Power Model
-To determine the operational carbon emissions from powering servers, we define the power usage $P^{i}=\bigl{(}p^{i}_{m,q}\bigl{)}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$.
-We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
-[…]
-(6) $p^{i}_{m,q}=\text{PUE}\cdot\left(p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(\text{util}^{i}_{q}\right)^{n}\right),$
-where
-(7) $\text{util}^{i}_{q}=\frac{\sum_{u\in\mathcal{U}}a^{i}_{u,q}}{\sum_{m\in\mathcal{M}}\left(d^{i}_{m,q}\cdot k_{m,q}\right)}.$
-Cloud tenant perspective.
-Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions.
-Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change.
-Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor.
-We assume that cloud providers attribute a constant, load-independent factor $P_{m}^{\text{attr}}$ to each active machine $m$:
-(8) $p^{i}_{m,q}=p_{m}^{\text{attr}}$
-3.4. Carbon Emission Model
-We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ with respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2):
-(9) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\cdot\left(p^{i}_{m,q}\cdot C^{i}+C_{m}^{\text{emb}}\right)$
-Depending on methodology, embodied emissions might not always be reported with respect to infrastructure usage, in which case $C_{m}^{\text{emb}}=0$.
-3.5. Optimization Problem
-We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies $\text{QoR}_{\text{target}}$:
+Carbon emission model.
+We define a service’s estimated carbon emissions $E^{i}$ during time interval $i$ with respect to the deployments current power usage, the region’s carbon intensity $C^{i}$ (in gCO2/kWh), and the attributed embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2) for running machine $m$ for time $\Delta$:
+(2) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\ \left(\Delta p^{i}_{m,q}\ C^{i}+C_{m}^{\text{emb}}\right)$
+Depending on the methodology, embodied emissions may be excluded from scheduling decisions (Bashir et al., 2024b), in which case $C_{m}^{\text{emb}}=0$.
+We do not explicitly model emissions per user or request in this paper, but they can be derived using carbon attribution models (Han et al., 2024; Westerhof et al., 2023).
+Optimization problem.
+We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies a $\text{QoR}_{\text{target}}\in[0,1]$:
 (3) $\min\quad$ $\sum_{i=\alpha}^{\omega}E^{i}$
@@
 (5) $\sum_{u\in\mathcal{U}}a^{i}_{u,q}\leq\sum_{m\in\mathcal{M}}d^{i}_{m,q}\cdot k_{m,q}$ $\forall i,\forall q$
-[…]
 (6) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
-where
-  • All requests served (11) ensures that all requests during each interval $\mathcal{T}_{i}$ are attributed to a quality tier.
-  • Sufficient resources (12) ensures that the provided machines have sufficient capacity to serve all requests per quality tier.
-  • Respect SLOs (13) ensures that all SLOs are respected.
-  • Sufficient QoR (14) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
-Note, that this model assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$) and can therefore not be solved in practice.
-However, we use it for assessing upper-bound potential and as the foundation of the online approach proposed in Section 4.
-3.6. Complexity
-The proposed optimization problem is an NP-hard mixed-integer linear program (MILP), as we minimize a linear objective function subject to linear constraints, which include discrete decision variables in $D^{i}$.
-Equivalence of power models.
-From the perspective of data center operators, the load-dependant power model from (6) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$.
-This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve.
-However, during optimization, we can relax (6) to the same form as (8), as long as $n\leq 1$, i.e.
-the power model is concave.
-Theorem 3.1.
-We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$.
-When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
-$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
-can be simplified to
-$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
-as it always yields an equivalent optimal deployment $D^{i}$.
-The proof is provided in Appendix A.
-Intuitively, only if the power-law relationship is strictly convex ($n>1$), it can be more efficient to deploy multiple machines at lower utilization rather than fully utilizing fewer machines.
-However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007).
-Hence, for $n\leq 1$, we can use a load-independent power model during optimization, as it yields the same optimal $D^{i}$.
+where (4) ensures that all requests during each interval $i$ are attributed to a quality tier, (5) ensures that the provided machines have sufficient capacity to serve all requests per quality tier, and (6) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
+Note, that this problem is not only NP-hard (see Appendix B) but also assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$). As such, it cannot be solved directly in practice.
 3. Online Approach
-For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$.
-Its performance is limited by two factors:
+A practical approach for carbon-aware QoR adaptation must cope with two sources of inaccuracy:
   1. (1)
-    the error of request and carbon intensity forecasts
+    Approximation errors: Optimally solving the MILP over a large solution space is computationally infeasible.
+In our experiments, no annual-horizon instance solved within 1 hour, while the median runtime on daily-horizon instances was 1.2 seconds (see Appendix C).
   2. (2)
-    the error introduced by approximating the NP-hard optimization problem
-In our experiments, conducted with realistic forecasts and strict time limits for solving the MILP, we achieve emission savings of $82\pm 6$ relative to the upper-bound potential (Section 5.4).
-4.1. Multi-Horizon Optimization
-Based on the optimization problem introduced in Section 3.5, we propose a forecast-based multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
-Forecasting.
-There exist various approaches for long-term load forecasting, such as Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011).
-As the electricity mix of regions does not change abruptly, long-term forecasts for ACI can be usually estimated reasonably well by fitting daily, weekly, and annual seasonalities.
-However, especially in regions that are heavily dependent on weather-dependent renewable power production, additionally using short-term forecasts can provide significantly higher accuracy (Maji et al., 2023; Li et al., 2024c).
-For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
-Algorithm.
-The approach is described in Algorithm 1.
-As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality.
-To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
+    Forecast errors: Long-term forecasts tend to be less accurate than short-term ones, particularly in regions with volatile renewable generation (Maji et al., 2023; Li et al., 2024b).
+Based on these findings, we propose a multi-horizon optimization which periodically updates forecasts and determines an optimal (or close-to-optimal) deployment and load balancing for every interval $i$, described in Algorithm 1.
+We perform the optimization in two distinct steps:
+  1. (1)
+    A long-term optimization, ensures global feasibility, i.e., meeting QoR constraints.
+It is executed every $\tau$ intervals (e.g., every 24 hours) and solves the MILP for the remainder of the year.
+Any past $D^{i}$ and $A^{i}$ are fixed, so we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval.
+Solutions may be approximate due to solver time limits.
+  2. (2)
+    A short-term optimization adapts to recent forecast updates and improves suboptimal long-term decisions.
+It is executed every interval and solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization).
+In Line 7 we assume $\gamma$ as the horizon..
+As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization has a significantly smaller search space and is more likely to find an optimal solution quickly.
+If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies Equation (5).
+After optimization, the determined deployment $D^{\alpha}$ is provisioned for the current interval.
+As short-term load and carbon intensity predictions are usually very precise, and the provisioning of new nodes can take significant time (Microsoft reports 6-8 minutes for creating a new LLM inference instance (Stojkovic et al., 2024)) we assume that there are usually no rapid auto-scaling decisions within an interval.
+After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
 Algorithm 1 Multi-Horizon Optimization
@@
 10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
-  1. (1)
-    A long-term optimization, executed every $\tau$ intervals (for example, every 24 hours), solves the MILP for the remainder of the year.
-Any $D^{i}$ and $A^{i}$ in the past are fixed, i.e.
-we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval.
-In this step, longer execution times and less optimal solutions are acceptable.
-  2. (2)
-    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization).
-In Line 7 we assume $\gamma$ as the horizon..
-As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly.
-If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (12) constraint.
-After optimization, the determined deployment $D^{\alpha}$ is provisioned for the current interval.
-As short-term load and carbon intensity predictions are usually very precise, and the provisioning of new nodes can take significant time (Microsoft reports 6-8 minutes for creating a new LLM inference instance (Stojkovic et al., 2024)) we assume that there are usually no rapid auto-scaling decisions within an interval.
-After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
-4.2. Automatic QoR Adaptation
-The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$.
-However, especially for best-effort users, another relevant question is: What are the expected annual emissions from this user group, and can we keep them within a predefined budget?
-In Section 5.4, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected annual emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year.
-Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:
-(15) $\max\quad$ $\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
-[…]
-(16) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
-In the multi-horizon optimization (Algorithm 1), we replace the long-term optimization in Line 5 with this objective:
-(17) $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\max_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\left(\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=0}^{I-1}\right)$
-Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate emissions over all intervals in (16).
-This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$.
-Past values in $D^{i}$ and $A^{i}$ are fixed.
 4. Experiments
-We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
-We use Gurobi (Gurobi Optimization, LLC, 2024) to solve optimization problems. All simulations were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job.
-Table 2.
-Upper-bound potential and carbon savings under realistic conditions in % for all request traces and regions at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$.
-Note, that other $\text{QoR}_{\text{target}}$ can lead to even better results (Figure X).
-Region Artificial Real Synthetic Mean
-Static Random Wiki (en) Wiki (de) Taxi Cell B Cell D Cell F
-NL 8.2 / 7.4 8.5 / 6.3 7.6 / 6.5 10.2 / 9.3 11.3 / 8.7 7.8 / 6.7 7.6 / 6.6 8.0 / 7.0 8.7±1.3 / 7.3±1.0
-CISO 7.0 / 6.8 7.5 / 6.0 7.5 / 7.0 10.0 / 9.4 11.3 / 9.2 6.6 / 6.1 6.3 / 5.9 6.8 / 6.4 7.9±1.7 / 7.1±1.3
-ES 7.1 / 6.3 7.3 / 5.2 6.5 / 5.6 9.4 / 8.4 10.4 / 8.0 6.6 / 5.6 6.6 / 5.7 6.9 / 5.9 7.6±1.4 / 6.3±1.1
-AU-QLD 7.9 / 7.8 8.4 / 7.0 7.1 / 6.5 6.1 / 5.8 8.8 / 6.9 7.6 / 7.0 6.9 / 6.4 8.0 / 7.5 7.6±0.8 / 6.9±0.6
-DE 6.3 / 5.4 6.7 / 4.5 5.7 / 4.6 8.7 / 7.6 9.7 / 6.9 5.9 / 4.8 5.8 / 4.8 6.1 / 5.1 6.9±1.4 / 5.5±1.1
-PL 4.1 / 3.8 4.5 / 2.9 3.5 / 2.9 6.7 / 6.2 7.5 / 5.7 3.8 / 3.3 3.7 / 3.2 4.1 / 3.5 4.7±1.4 / 3.9±1.2
-ERCOT 4.2 / 3.1 4.6 / 2.6 3.6 / 2.4 6.3 / 5.1 7.4 / 5.8 3.9 / 2.8 3.7 / 2.5 4.1 / 3.0 4.7±1.3 / 3.4±1.2
-SE 2.7 / 2.4 3.0 / 1.6 2.0 / 1.5 4.9 / 4.5 6.1 / 5.4 2.3 / 1.9 2.2 / 1.7 2.6 / 2.2 3.2±1.4 / 2.6±1.4
-NYISO 2.3 / 1.8 2.7 / 1.1 1.7 / 1.1 4.6 / 4.3 5.6 / 4.6 1.9 / 1.5 1.8 / 1.3 2.2 / 1.7 2.8±1.4 / 2.2±1.3
-PJM 1.8 / 1.7 2.3 / 0.8 1.3 / 1.0 4.1 / 4.0 5.5 / 4.3 1.6 / 1.4 1.4 / 1.1 1.8 / 1.6 2.5±1.4 / 2.0±1.3
-5.1. Experimental Setup
-As ACI is usually reported hourly, we define $\Delta=1$ hour.
-[…]
+We evaluate our approach by simulating an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
+All experiments were provided 16 cores of an HPC cluster, using Gurobi (Gurobi Optimization, LLC, 2024) for optimization.
 […]
 Request traces.
-For our analysis, we require 4 years of hourly request traces: 2020-2022 for fitting forecasting models and 2023 for the analysis itself.
-As such data are not publicly available for LLM inference servies, we show the effectiveness of our method across a diverse set of other real and synthetic request traces:
+We require 4 years of hourly request traces: 3 for fitting forecasting models and 1 for the analysis itself.
+As such data are not publicly available for LLM inference services, we evaluate our method across a diverse set of other request traces:
   • Artificial: A Static trace that assumes a constant stream of hourly requests, and a Random trace, where we sample hourly requests from a normal distribution.
-  • Real: Two real-life traces representing the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n.
-d.]), Wiki (en) and Wiki (de), and a trace based on the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
-  • Synthetic: Three generated traces, called Cell B, D, and F, which are based on instance event traces of Borg cluster cells at Google (Google, 2019).
-All request trace datasets and our forecast methodology are explained in detail in Appendix B.
-Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
-  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
-  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
-  • In Australia, we consider Queensland (AU-QLD)
+  • Human-Generated: Two real-world traces from unrelated domains, namely the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n.
+d.]), Wiki (en) and Wiki (de), and the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
+  • Synthetic: Three traces generated from models fitted to Google’s Borg cluster cells (Google, 2019), called Cell B, D, and F.
+All request trace datasets and our forecast methodology are explained in detail in Appendix D.
+Carbon intensity traces.
+We evaluate on ten globally distributed regions: California, US (CISO), Texas, US (ERCOT), New York, US (NYISO), Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), Sweden (SE), and Queensland, Australia (AU-QLD).
 All carbon intensity data was provided by ElectricityMaps (Maps, 2024).
-We normalized all time zones to align the time of day with the request traces.
-The temporal variability of regions is depicted in Figure 1.
-Our forecast methodology is explained in Appendix C.
+Our forecast methodology is explained in Appendix E.
 4.1. Absolute Emissions
-We first analyze the absolute annual emissions without carbon-aware QoR adaption, i.e.
-for $\gamma=1$ hour.
-The results reflect emission savings driven purely by increased energy demand from serving more requests on Llama 3.1 70B ($\text{QoR}_{\text{target}}\to 1$).
-Figure 4 shows the annual emissions for Wiki (en)—the trend is equivalent for all datasets.
-We observe very large differences across different regions: For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally.
-Serving all requests with the 70B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source.
-As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$.
-Figure 4. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
-Figure 5. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains.
+We first assess the absolute annual emissions without carbon-aware QoR adaption.
+Figure 3 shows the annual emissions for Wiki (de).
+As expected, emissions scale linearly with $\text{QoR}_{\text{target}}$, but we observe large differences across regions.
+For example, operating the service in Sweden emits almost 27 times less carbon than in Poland.
+For reference, ChatGPT, one of the biggest LLM services today, reached over 200 million active users in 2024 (Reuters, 2024).
+Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour—two orders of magnitude more than our Wiki (en) trace.
+Assuming similar performance as Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers already exceed multiple 10,000 tCO2.
+Figure 3. Annual operational carbon emissions for Wiki (de) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
 4.2. Upper-Bound Potential
 […]
-Carbon savings beyond energy efficiency.
-In blue (left), Table 2 presents the relative additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week compared to $\gamma=1$ hour, where emissions directly depend on energy demand.
-We observe that through carbon-aware QoR adaptation, emissions drop by around 8 % in many regions.
-While request trace patterns exhibit little difference in potential, traces with a low volume of requests (Wiki (de) and Taxi) show increased potential, due to the higher impact of discrete scheduling decisions.
+Table 1 presents the potential for additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week.
+That is, instead of meeting the 50 % target hourly, we ensure that 50 % of requests are served by LLaMA 3.1 70B over each week.
+We observe that this increased flexibility enables emissions savings of around 8 % in many regions.
+While request trace patterns exhibit little difference in potential, low-volume traces (Wiki (de) and Taxi) offer greater savings due to the larger impact of discrete scheduling decisions.
 For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines.
-In experiments with even fewer machines we managed to gain relative savings of more than 40 %—however, we do not analyze such cases further, as we focus on high-impact scenarios with many machines.
-Impact of validity periods.
-Figure 5 reports the relative and absolute savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods.
+Note, that the potential depends heavily on $\text{QoR}_{\text{target}}$ as shown in Appendix F.
+Table 1. Relative upper-bound carbon savings potential.
+Region Artificial Human-Generated Synthetic Mean
+Static Rnd. Wiki Wiki Taxi Cell Cell Cell
+(en) (de) B D F
+NL 8.2 8.5 7.6 10.2 11.3 7.8 7.6 8.0 8.7±1.3
+CISO 7.0 7.5 7.5 10.0 11.3 6.6 6.3 6.8 7.9±1.7
+ES 7.1 7.3 6.5 9.4 10.4 6.6 6.6 6.9 7.6±1.4
+AU-QLD 7.9 8.4 7.1 6.1 8.8 7.6 6.9 8.0 7.6±0.8
+DE 6.3 6.7 5.7 8.7 9.7 5.9 5.8 6.1 6.9±1.4
+PL 4.1 4.5 3.5 6.7 7.5 3.8 3.7 4.1 4.7±1.4
+ERCOT 4.2 4.6 3.6 6.3 7.4 3.9 3.7 4.1 4.7±1.3
+SE 2.7 3.0 2.0 4.9 6.1 2.3 2.2 2.6 3.2±1.4
+NYISO 2.3 2.7 1.7 4.6 5.6 1.9 1.8 2.2 2.8±1.4
+PJM 1.8 2.3 1.3 4.1 5.5 1.6 1.4 1.8 2.5±1.4
+Figure 4.
+Relative (top) and absolute (bottom) additional carbon savings for different validity periods of our proposed online approach (colored) compared to upper bound potential (gray).
+Larger validity periods enable higher savings.
+Figure 4 reports the relative and absolute saving potential for Wiki (de) at $\text{QoR}_{\text{target}}=0.5$ and different validity periods in gray.
 Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings.
@@
 As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements.
-In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity.
-Impact of premium users.
-Figure 6 illustrates the loss of potential with a growing fraction of premium users at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$.
-We observe a slightly sub-linear decrease with a certain offset which depends on the overall request volume and, hence, the impact of discrete scheduling decisions.
-For reference, ChatGPT reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024).
-The resulting 5 % premium users decrease the potential savings by 5.6 % on average.
-Figure 6. Impact of premium users on the potential savings.
-Impact of $\text{QoR}_{\text{target}}$.
-Figure 7 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$.
-At $\text{QoR}_{\text{target}}=0$, we serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings.
-Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B.
-A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
-Figure 7. Additional relative savings for different $\text{QoR}_{\text{target}}$.
-Figure 8. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?
-Periods of low service quality.
-While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 8.
-For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right).
-However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left).
-We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR.
-Consequently, very large validity periods of more than a month are not practical in real-world scenarios.
-5.4. Evaluation under Realistic Conditions
-We evaluated the practicability of carbon-aware QoR adaptation using the proposed multi-horizon optimization under real forecasts (see Appendix B as well as C) and approximated MILP solutions: Long-term optimizations are stopped at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 30s; short-term optimizations at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 10s.
-[…]
-Figure 9. Additional carbon savings under realistic conditions for Wiki (de) across all regions and $\text{QoR}_{\text{target}}=0.5$.
-Outlook of annual emissions.
-Long-term optimizations are computationally cheap when accepting near-optimal solutions—within just 30 seconds, the mean $\text{MILP}_{\text{gap}}$ reached 0.53 % accross all experiments.
-This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts, to represent, for instance, worst-case or best-case scenarios.
-Figure 10 demonstrates emission outlooks for Wiki (en) traces in California.
-For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively.
-Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
-(a) Annual emission outlook at three discrete points in time.
-(b) Annual emission outlook; weekly for $\text{QoR}_{\text{target}}=0.5$.
-Figure 10. Projected emissions (blue) based on long-term forecasts. The interval shows the range between best- and worst-case forecasts. Red indicates emissions already incurred.
-Figure 10(a) shows the resulting outlooks a the start of the year, after 10 weeks, and after 40 weeks.
-The outlooks provide a comprehensive picture of the future, based on the currently available information.
-From the initial outlook (left), service providers may conclude that setting $\text{QoR}_{\text{target}}=0.5$ is a reasonable target to stay within an annual carbon budget of 1000 tCO2.
-Consequently, we optimize for $\text{QoR}_{\text{target}}=0.5$ for the remainder of the year.
-Figure 10(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$.
-The upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as incurred emissions accumulate.
-5.5. Automatic QoR Adaptation
-Figure 11. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL). Our approach provides the most consistent QoR throughout the year.
-Lastly, we illustrate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.2 on Wiki (en) in three exemplary regions.
-For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment (Section 5.3) for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$.
-Consequently, under perfect forecasts, the optimal approach would be to choose $\text{QoR}_{\text{target}}=0.5$ for every time step.
-We compare our approach to two greedy baselines that have access to the same forecasts as our approach:
-  1. (1)
-    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves (15). Any remaining or overspent budget is evenly distributed across the remaining time intervals.
-  2. (2)
-    Greedy (weighted) also solves (15), but weights the hourly budgets based on forecasted load and carbon intensity.
-Every 24 hours, as request and carbon intensity forecasts are updated, weightings get adjusted.
-Figure 11 shows the daily QoR of the greedy baselines and our automatic QoR adaptation.
-Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year.
-Greedy (weighted) delivers more consistent QoR in all experiments but tends to fall back to either $\text{QoR}=1$ (NL) or $\text{QoR}=0$ (CISO and NYISO) toward the end of the year due to forecast errors.
-In the CISO and NYISO regions, it even exceeds the budget by 0.8 and 1.2 %, respectively.
-Our approach provides the most consistent service across all experiments as it periodically adjusts $\text{QoR}_{\text{target}}$ based on long-term optimizations.
-The annual budget is utilized to 98.2–99.8 %.
-6. Discussion
-This section highlights the impact of our results on energy-intensive services and discusses the limitations of this study.
-Impact.
-We emphasize the significant potential for carbon efficiency improvements in cloud services, based on our analysis in Section 5.2.
-ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024).
-Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our larger scenarios.
-Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
-Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy.
-However, and increasing part of the community is advocating to base load shifting decisions based on marginal signals (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids.
-Yet, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone.
-Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion.
-We predict that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods.
-Our approach is directly applicable to MCI signals or any other scalar metric without requiring adaptations.
-Limitations and future work.
-This paper presents the first analysis on the potential of carbon-aware QoR adaptation in services, focusing on an LLM inference use case.
-While our approach is already applicable to more complex scenarios with multiple service quality tiers and heterogeneous infrastructure, such cases have not yet been investigated experimentally.
-Second, our current modeling only considers the QoR for user groups, rather than individual users.
-This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven.
-For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group.
-We consider fairness aspects with more granular user modeling as an area for future work.
-Third, future work should explore the impact of adaptive QoR adjustments on user quality of experience and the resulting implications on user behavior.
-For instance, in the context of generative AI, reduced QoR might incentivize users to resubmit requests, potentially offsetting carbon savings.
-Unfortunately, to the best of our knowledge, there are no established behavioral models for generative AI usage.
-Developing such models would be valuable to better quantify the real-life impacts of sustainability approaches, including rebound effects.
-7. Related Work
-We survey related works on carbon-aware computing as well as carbon efficiency in energy-intensive cloud services.
-Carbon-aware cloud services.
-The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy.
-Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a; Lechowicz et al., 2024).
-However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
-Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed data centers and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016).
-For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint under latency constraints.
-CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks.
-Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity.
-However, there are no existing approaches that address services which must operate in a single region due to data, security, or infrastructure constraints.
-Sustainability in generative AI inference.
-Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023; Samsi et al., 2023).
-Especially, the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024).
-Many recent approaches target sustainability aspects of inference services throuh power-aware scheduling (Stojkovic et al., 2024, 2025; Qiu et al., 2025; Patel et al., 2024) or by considering their carbon and water footprint in geographic load balancing (Li et al., 2024d).
-Service quality adaptation.
-While service degradation is a common strategy to prevent system overload (Meza et al., 2023; Park et al., 2024), recent research has proposed designing applications with different service quality tiers as a means of improving sustainability (Vitali, 2022).
-In the context of carbon-aware LLMs, we identified only one approach, called Sprout (Li et al., 2024b), which optimizes the autoregressive generation process.
-Sprout returns shorter responses during periods of high carbon intensity, as the energy demand directly correlates with the number of generated tokens.
-Thereby, it effectively adapts QoR over time and can be further enhanced by leveraging the optimized infrastructure deployments and load balancing proposed in this paper.
+In contrast, regions like Sweden, New York, and PJM exhibit little temporal variation in carbon intensity and have therefore a low savings potential of below 5 %.
+4.3. Potential under Realistic Conditions
+We evaluate the practicability of carbon-aware QoR adaptation by applying our online approach under real forecasts (see Appendix D and E) and approximated MILP solutions: Long-term optimizations are stopped after 30s; short-term optimizations after 10s.
+Figure 4 shows the gap to the upper-bound performance.
+Across all experiments, savings under realistic conditions reached $82\pm 6$ of the upper-bound potential.
+Notably, even in regions with highly unpredictable request patterns (Random and Cell B, D and F), the online optimization does not show a significant performance drop.
 5. Conclusion
-As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets.
-In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency.
-Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand.
-Moreover, we propose an approach to automatically adapt the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
+In this paper, we presented a first analysis of carbon-aware QoR adaptation, a strategy to reduce operational carbon emissions of cloud services without relying on geo-distributed load balancing or delay-tolerant workloads.
+Our results demonstrate that adjusting service quality in response to carbon intensity can reduce emissions by up to 10 % beyond savings related to increased energy efficiency.
+To cover a wider range of realistic settings, future work should extend our analysis to scenarios with heterogeneous machines, multiple quality tiers, and diverse user groups.
+It is also important to examine user-side effects, such as whether low QoR might trigger behavioral responses like repeated requests, potentially offsetting carbon savings.
+Additionally, future work should explore the potential of QoR adaptation under sustainability metrics used for short-term demand management, like marginal carbon intensity or excess power during curtailments (Wiesner and Kao, 2025).
+These metrics exhibit higher temporal variability than average carbon intensity, potentially enabling higher savings and the use of shorter validity periods.
 Acknowledgements.
-This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. 01IS18025A and 01IS18037A).
-Appendix A Proof of Theorem 3.1
-Theorem 3.1 We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$.
-When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
-$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
-can be simplified to
-$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
-as it always yields an equivalent optimal deployment $D^{i}$.
-Lemma A.1.
-Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines than strictly necessary.
-Formally, if $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$, an utilization-dependent power model $f$ is strictly subadditive
-(18) $f(x+y)<f(x)+f(y),$
-where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
-Proof of Lemma A.1.
-We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
-Case 1: Linear power model ($n=1$):
-(19) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
-$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
-(20) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
-(21) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
-(22) $f(x+y)$ $<f(x)+f(y)$
-Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $f(0)\geq 0$, $f$ is strictly subadditive on $(0,\infty)$.
-(23) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
-(24) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
-(25) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl{(}}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl{(}}}_{>1}<0$
-(26) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
-Additionally, $f(0)=p_{m,q}^{\text{max}}>0$. ∎
-Proof of Theorem 3.1.
-For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
-[…]
-  1. (1)
-    Feasibility with fewer machines is impossible. If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate Sufficient resources (12).
-  2. (2)
-    Using extra machines is suboptimal.
-By Lemma A.1, under the utilization-dependent power model, there is no incentive to deploy extra machines.
-Likewise, any extra machine increases the objective under the simplified power model.
-We conclude that, for both power models, any optimal solution must yield an equivalent deployment $\check{D}^{i}$ for each interval $i$. ∎
-Appendix B Request Traces
-This section describes all request trace datasets used in the experiments. The traces are visualized in Figure 12; Table 3 presents summary statistics.
+This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref.
+BIFOLD24B), the German Research Foundation (DFG) as FONDA—Foundations of Workflows for Large-Scale Scientific Data Analysis (ref.
+CRC 1404), and the EU Horizon project EXIGENCE under Grant Agreement No.
+101139120.
+Appendix A Nomenclature
+Table 2. Description of symbols. Small letters denote the elements of the corresponding, capitalized matrix.
+Input variables
+$\mathcal{U}$ Set of user groups $u$
+$\mathcal{M}$ Set of machine types $m$
+$\mathcal{Q}$ Tuple of two service quality tiers $q$
+$\gamma\in\mathbb{N}$ Validity period length
+$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $i$.
+$C^{i}\in\mathbb{R}$ Carbon intensity during $i$.
+$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $i$.
+$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per interval $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
+$C_{m}^{\text{emb}}\in\mathbb{R}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
+Decision variables
+$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $i$.
+$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $i$.
+Appendix B Proof of NP-Hardness
+We prove that the optimization problem defined in Equations (3)–(6) is NP-hard via a polynomial-time reduction from the Bin Packing problem, which is known to be NP-complete (Garey and Johnson, 1979).
+Bin Packing is defined as: Given a set of items with sizes $a_{j}\in(0,1]$ and an integer $B$, can the items be packed into at most $B$ bins of capacity 1?
+Consider a simplified instance of our problem with a single interval $i$, one user group $u$, one machine type $m$, and one quality tier $q$.
+Each machine has capacity $k_{m,q}=1$.
+Requests correspond exactly to Bin Packing items, each with size $a_{j}$, and the total request volume is $\sum_{j}a_{j}$.
+In this setting, determining feasibility under Constraints (4) and (5) is equivalent to solving the Bin Packing problem: each machine corresponds to a bin, each request to an item, and the capacity constraint ensures that the sum of assigned request sizes per machine does not exceed 1.
+Since this special case of our problem is equivalent to Bin Packing, and Bin Packing is NP-complete, it follows that our optimization problem is NP-hard.
+$\square$
+Appendix C Performance Analysis
+To illustrate the performance improvements achieved by the proposed multi-horizon optimization, we analyze 240 experiments (all regions, all trace datasets, and $\text{QoR}_{\text{target}}\in\{0.3,0.5,0.7\}$) for the annual-horizon and daily-horizon optimization.
+These correspond to the long-term and short-term optimization components of our evaluated online approach.
+Figure 5. CDF of optimization gap over time.
+Figure 5 shows a CDF over the fraction of runs below a certain optimization gap over time.
+In the annual-horizon optimization, around 10 % of all runs have not found a solution within 1 % of the optimum within 10 minutes, while no run found an optimal solution within one hour.
+In the daily-horizon optimization, the median runtime was only 1.2 seconds and more than 80 % of all runs found an optimal solution within 10 seconds.
+Appendix D Request Traces
+This section describes all request trace datasets used in the experiments. Table 3 presents summary statistics.
 Table 3. Request trace and 24-hour forecast statistics.
@@
 Cell F 1.58 ± 0.41 0.87 4.32 18.2 ± 9.3
-Figure 12. First month of data and KDE over all data points in 2023 for the eight datasets used in the evaluation.
 Artificial Datasets.  We artificially create two traces:
@@
 We apply the same forecasting methodology that is used for real datasets.
-Appendix C Carbon Intensity Forecasts
+Appendix E Carbon Intensity Forecasts
 For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets.
@@
 Short-term forecasts are updated daily at midnight.
-Table 4. Carbon forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
+Table 4. Carbon intensit forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
 Region Day 1 Day 2 Day 3 Day 4
@@
 AU-QLD 3.93 3.98 4.06 5.87
+Appendix F Impact of QoRtarget
+Figure 6 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$ and validity periods, for Wiki (en) in Germany and California.
+At $\text{QoR}_{\text{target}}=0$ and $\text{QoR}_{\text{target}}=1$, we serve all requests using either LLaMA 3.1 8B or 70B, leaving no flexibility for additional savings.
+A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
+Figure 6. Additional relative savings for different $\text{QoR}_{\text{target}}$.
+Appendix G Periods of Low Service Quality
+While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 7.
+For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right).
+However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left).
+We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR.
+Consequently, large validity periods are not practical in many real-world scenarios.
+Figure 7. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?

### v4

@@
 In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers.
-We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency and introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings.
+We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency.
+Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget.
+Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
 Sustainable computing, quality of service, LLM inference, green AI
-journalyear: 2025copyright: ccconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computing
+journalyear: 2025copyright: noneconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingThis manuscript is an extended version of our paper published at ACM e-Energy’25 (DOI: 10.1145/3679240.3734614).
+Compared to the published version, we (1) add a time-based vs.
+utilization-based power attribution perspective together with a proof that both yield equivalent provisioning decisions under mild assumptions (Appendix H), (2) extend the online approach with an automatic adaptation of $\text{QoR}_{\text{target}}$ to meet a fixed annual carbon budget (Sections 3.1 and 4.4).
 1. Introduction
@@
 We examine the impact of dynamically adjusting service QoR with respect to temporal changes in ACI to enhance carbon efficiency without relying on geo-distributed load balancing.
-[…]
+Problem setting.
+Given an interval duration $\Delta$, we divide our optimization time window $[0,T]$ into $I=T/\Delta$ time intervals, indexed by $i$.
+For the remainder of this paper, we consider $T=1~\text{year}$ and $\Delta=1$ hour.
+We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$, and a tuple of two service quality tiers $\mathcal{Q}=(\text{Tier~1},\text{Tier~2})$.
+Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
+For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group.
+Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$.
+The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$).
+The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval.
+Appendix A shows an overview of all input and decision variables.
 Quality of Responses.
@@
 We do not explicitly model emissions per user or request in this paper, but they can be derived using carbon attribution models (Han et al., 2024; Westerhof et al., 2023).
+Time-based vs.
+utilization-based power attribution.
+The operational emissions in (2) require a choice of how to attribute power draw to a service and its resources.
+Two common modeling choices are: (i) utilization-based power attribution, where per-machine power depends on resource utilization, and (ii) time-based power attribution, where every active machine is assigned a (typically constant) attributed power draw per unit time.
+Utilization-based models are a natural fit when modeling operator-side electricity usage (often reported as Scope 2), where metering reflects the actual energy consumed by servers.
+Time-based attribution is common in tenant-side carbon accounting (often reported as Scope 3) when emissions are allocated proportional to instance-hours.
+However, Scope 3 methodologies can also be utilization- or energy-based, and Scope 2 reporting can rely on attributed values depending on available measurements.
+For utilization-based attribution, machine power draw can be modeled as utilization-dependent, e.g., with a concave power law,
+(3) $p^{i}_{q}=p^{\text{idle}}+\bigl(p_{q}^{\text{max}}-p^{\text{idle}}\bigr)\cdot\bigl(\text{util}^{i}_{q}\bigr)^{n},\quad n\in(0,1],$
+where, in our setting with one user group and one machine type, $\text{util}^{i}_{q}=a^{i}_{q}/(d^{i}_{q}\cdot k_{q})$.
+Under time-based attribution, $p^{i}_{q}$ is treated as utilization-independent (constant per active instance).
+We show that under mild concavity assumptions, optimizing under time-based vs.
+utilization-based attribution yields equivalent provisioning decisions (Appendix H).
 Optimization problem.
@@
 Based on these findings, we propose a multi-horizon optimization which periodically updates forecasts and determines an optimal (or close-to-optimal) deployment and load balancing for every interval $i$, described in Algorithm 1.
+Algorithm 1 Multi-Horizon Optimization
+1:for $\alpha\leftarrow 0$ to I1I-1 do
+2: # Long-term optimization
+3:if $\alpha\equiv 0\pmod{\tau}$ then
+4:  $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
+5:  $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\sum_{i=\alpha}^{I-1}E^{i}$
+6: # Short-term optimization
+[…]
+8: # Progress in time
+[…]
+[…]
 We perform the optimization in two distinct steps:
@@
 After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
-Algorithm 1 Multi-Horizon Optimization
-1:for $\alpha\leftarrow 0$ to $I-1$ do
-2:    # Long-term optimization
-3:   if $\alpha\equiv 0\pmod{\tau}$ then
-4:      $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
-5:      $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\sum_{i=\alpha}^{I-1}E^{i}$
-6:    # Short-term optimization
-7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}}\sum_{i=\alpha}^{I-1}E^{i}$
-8:    # Progress in time
+3.1. Automatic QoR Adaptation under an Annual Budget
+The optimization problem in Section 3 minimizes emissions under a given $\text{QoR}_{\text{target}}$.
+However, especially for best-effort users, an alternative objective is to provide the highest possible service quality while staying within an annual carbon budget $B$.
+This setting is relevant when operators (or tenants) face internal sustainability targets, reporting requirements, or carbon pricing that effectively turns emissions into a constrained resource.
+We model this by making $\text{QoR}_{\text{target}}$ a decision variable and optimizing
+(8) $\max\quad$ $\text{QoR}_{\text{target}}$
 […]
-10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
+(9) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
+(10) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}.$
+In practice, when re-optimizing at time $\alpha$, a part of the budget has already been spent.
+Let $E^{0},\dots,E^{\alpha-1}$ be the (observed) emissions of past intervals.
+Then the remaining budget is
+(11) $B_{\text{rem}}(\alpha)=B-\sum_{i=0}^{\alpha-1}E^{i},$
+and the long-term optimization step is solved over the remaining horizon with $\sum_{i=\alpha}^{\omega}E^{i}\leq B_{\text{rem}}(\alpha)$ while keeping all past decisions fixed.
+In the multi-horizon optimization (Algorithm 1), we use (8) as the objective for the long-term optimization step.
+This enables periodic updates of $\text{QoR}_{\text{target}}$ based on forecast updates and already-incurred emissions, while the short-term step continues to refine per-interval decisions within a smaller horizon.
 4. Experiments
@@
 All experiments were provided 16 cores of an HPC cluster, using Gurobi (Gurobi Optimization, LLC, 2024) for optimization.
-[…]
+Scenario.
+We investigate an LLM inference scenario, where a single user group is served by a LLaMA 3.1 (Dubey et al., 2024) model through the inference and serving engine vLLM (Kwon et al., 2023).
+Users can be fully served by LLaMA 3.1 8B ($\text{QoR}=0$), LLaMA 3.1 70B ($\text{QoR}=1$), or a fraction of both models ($0<\text{QoR}<1$).
+We consider $\mathcal{M}=\{\text{EC2 p4d.24xlarge}\}$, currently the only instance type with high-performance GPUs available on AWS, and define $p_{m}^{\text{attr}}=3781.8\,$W and $C_{m}^{\text{emb}}=135.3$ gCO2 based on (Teads Engineering, 2024).
+We model the machine performance $K$ using recent benchmarks of vLLM on a EC2 p4d.24xlarge for inference, which state a throughput of 11.57 requests per second for LLaMA 3.1 8B and 5.05 requests per second for LLaMA 3.1 70B (vLLM Team, [n.
+d.]).
 Request traces.
@@
 All request trace datasets and our forecast methodology are explained in detail in Appendix D.
-Carbon intensity traces.
-We evaluate on ten globally distributed regions: California, US (CISO), Texas, US (ERCOT), New York, US (NYISO), Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), Sweden (SE), and Queensland, Australia (AU-QLD).
+Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
+  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
+  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
+  • In Australia, we consider Queensland (AU-QLD)
 All carbon intensity data was provided by ElectricityMaps (Maps, 2024).
+We normalized all time zones to align the time of day with the request traces.
 Our forecast methodology is explained in Appendix E.
@@
 4.2. Upper-Bound Potential
-[…]
+To assess the potential of carbon-aware QoR adaptation, we assume perfect forecasts, and solve Equation (4) for $\gamma=\{$8 hours, 1 day, 1 week, 1 month}\}.
+Shorter periods show limited potential, as ACI is reported hourly and tends to vary only slowly over time.
+Note, that long validity periods can lead to reduced service quality over extended periods, see Appendix G.
+We optimize the objective until it is within 0.1 % of the optimal solution or after 1 hour.
 Table 1 presents the potential for additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week.
@@
 PJM 1.8 2.3 1.3 4.1 5.5 1.6 1.4 1.8 2.5±1.4
-Figure 4.
-Relative (top) and absolute (bottom) additional carbon savings for different validity periods of our proposed online approach (colored) compared to upper bound potential (gray).
-Larger validity periods enable higher savings.
 Figure 4 reports the relative and absolute saving potential for Wiki (de) at $\text{QoR}_{\text{target}}=0.5$ and different validity periods in gray.
@@
 4.3. Potential under Realistic Conditions
+Figure 4.
+Relative (top) and absolute (bottom) additional carbon savings for different validity periods of our proposed online approach (colored) compared to upper bound potential (gray).
+Larger validity periods enable higher savings.
 We evaluate the practicability of carbon-aware QoR adaptation by applying our online approach under real forecasts (see Appendix D and E) and approximated MILP solutions: Long-term optimizations are stopped after 30s; short-term optimizations after 10s.
 Figure 4 shows the gap to the upper-bound performance.
-Across all experiments, savings under realistic conditions reached $82\pm 6$ of the upper-bound potential.
+Across all experiments, savings under realistic conditions reached $82\pm 6$ % of the upper-bound potential.
 Notably, even in regions with highly unpredictable request patterns (Random and Cell B, D and F), the online optimization does not show a significant performance drop.
+4.4. Automatic QoR Adaptation under an Annual Budget
+Lastly, we evaluate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 3.1 on Wiki (en) in three exemplary regions.
+For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$ (Section 4.2).
+Consequently, under perfect forecasts, an optimal solution would maintain $\text{QoR}_{\text{target}}\approx 0.5$ throughout the year.
+Baselines.  We compare our approach to two greedy baselines that use the same forecasts as our method:
+  1. (1)
+    Greedy (constant) divides the remaining annual budget into equal per-interval budgets and greedily selects the highest feasible per-interval quality given the current forecast.
+Any remaining or overspent budget is evenly distributed across the remaining intervals.
+  2. (2)
+    Greedy (weighted) follows the same approach but weights the per-interval budgets by forecasted load and carbon intensity. Weightings are updated whenever forecasts are updated.
+Figure 5. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL) under an annual carbon budget. Automatic adaptation provides consistent QoR throughout the year.
+Figure 5 shows the resulting daily QoR under realistic forecasts.
+Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year.
+Greedy (weighted) is more stable but can still drift toward extreme QoR values toward the end of the year as forecast errors accumulate.
+In contrast, periodic long-term re-optimizations allow our approach to adjust $\text{QoR}_{\text{target}}$ over time, yielding more consistent QoR while staying within the annual budget.
 5. Conclusion
@@
 Appendix E Carbon Intensity Forecasts
-For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets.
-As in (Lechowicz et al., 2024), we generate synthetic short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region, see Table 4.
-Short-term forecasts are updated daily at midnight.
 Table 4. Carbon intensit forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
@@
 AU-QLD 3.93 3.98 4.06 5.87
+For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets.
+As in (Lechowicz et al., 2024), we generate synthetic short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region, see Table 4.
+Short-term forecasts are updated daily at midnight.
 Appendix F Impact of QoRtarget
@@
 Figure 8. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?
+Appendix H Equivalence of Provisioning Under Time- vs. Utilization-based Power Attribution
+We prove that the optimal provisioning decision is the same under (i) utilization-based power attribution (3), where per-machine power depends on resource utilization, and (ii) time-based power attribution, where each active machine is assigned a utilization-independent attributed power draw.
+This is important, as utilization-based power attribution introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$.
+This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve.
+Assumption 1.
+$p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$. This reflects the basic property that active machines consume more power under load than when idle or turned off.
+Assumption 2.
+$n\leq 1$, i.e.
+the power model is concave, which is a common assumption for computing hardware as servers run more efficiently under high load (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007).
+Assumption 3.
+$p_{m_{1}}^{\text{idle}}<p_{m_{2}}^{\text{idle}}\ \Rightarrow\ p_{m_{1},q}^{\text{max}}<p_{m_{2},q}^{\text{max}}\quad\forall m_{1},m_{2}\in\widehat{\mathcal{M}}$.
+If a quality tier can be deployed on different machine types $\widehat{\mathcal{M}}$, we assume a strict ordering of these machines based on their idle and max power usage.
+Theorem 1.
+Let Assumptions 12, and 3 hold. When minimizing (4) subject to (5) – (7), the utilization-based power attribution,
+$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl)\cdot\bigl(\text{util}^{i}_{q}\bigl)^{n},$
+can be simplified to the same form as the time-based attribution
+$\tilde{p}^{i}_{m,q}=p_{m,q}^{\text{max}},$
+as it always yields an equivalent optimal deployment $D^{i}$.
+Lemma 0.
+Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines of the same type than strictly necessary.
+Formally, if Assumptions 1 and 2 hold, utilization-based attribution $f$ is strictly subadditive
+(12) $f(x+y)<f(x)+f(y),$
+where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
+Proof of Lemma 2.
+We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
+Case 1: Linear power model ($n=1$):
+(13) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
+$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
+(14) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
+(15) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
+(16) $f(x+y)$ $<f(x)+f(y)$
+Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $(0)\geq 0, ff$ is strictly subadditive on $(0,\infty)$.
+(17) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
+(18) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
+(19) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl(}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl(}}_{>1}<0$
+(20) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
+Additionally, $f(0)=p_{m}^{\text{idle}}>0$. ∎
+Proof of Theorem 1.
+For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
+[…]
+We show, that both power attribution models yield $\check{D}^{i}$:
+  • Feasibility with fewer machines is impossible: If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate (6).
+  • Using extra machines is suboptimal for time-based power attribution: For $\tilde{p}^{i}_{m,q}$, any extra machine increases the objective by $p_{m,q}^{\text{max}}$.
+  • Using extra machines is suboptimal for utilization-based power attribution: By Lemma 2, for $p^{i}_{m,q}$ there is no incentive to deploy extra machines of the same type.
+If a quality tier $q$ can be served by multiple machine types $\widehat{\mathcal{M}}$, under Assumption 3 (a strict ordering of these machines based on their idle and max power usage), if a more efficient machine type $m_{1}$ is still able to carry extra load, using a less efficient type $m_{2}$ instead can only increase the objective.
+Consequently, also across heterogeneous machine types, there is no incentive to add extra machines.
+In either case (homogeneous or strictly ordered heterogeneous machines), any optimal solution must use exactly the minimal number of machines of each machine type.
+Because both the time-based and the utilization-based models thus select the same minimal deployment $\check{D}^{i}$, they yield equivalent optimal deployments.
+∎